### v1

# Learning the Mechanism of Catastrophic Forgetting: A Perspective from Gradient Similarity

Abstract
Catastrophic forgetting during knowledge injection severely undermines the continual learning capability of large language models (LLMs). While existing methods attempt to mitigate this issue, they often lack a foundational theoretical explanation. In this paper, we establish a theoretical framework from the perspective of gradient to explain catastrophic forgetting. First, we theoretically prove that strongly negative gradient similarity is the fundamental cause of catastrophic forgetting. Subsequently, gradient similarity is employed to identify two types of neurons during catastrophic forgetting: conflicting neurons, which induce forgetting and account for 50%–75% of all neurons, and collaborative neurons, which mitigate forgetting and account for 25%-50%. Ultimately, a novel knowledge injection method named Collaborative Neural Learning (CNL) is proposed. By freezing conflicting neurons and only updating collaborative neurons, CNL theoretically eliminates catastrophic forgetting under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set. Studies on five LLMs, four datasets, and four optimizers prove that CNL achieves zero forgetting under in-set experiments, and reduces the level of catastrophic forgetting by 59.1%–81.7% under out-of-set experiments.
1 Introduction
Large language models (LLMs) have shown exceptional knowledge capacity and achieved remarkable potential across a broad range of domains Achiam et al. (2023); Touvron et al. (2023). To continuously expand their knowledge breadth, various knowledge injection methods have been proposed Lewis et al. (2020); Wang et al. (2024); Ke et al. (2023). Although these methods enable LLMs to acquire new knowledge, they also cause the forgetting of previously mastered knowledge Zhou et al. (2025). This phenomenon, known as catastrophic forgetting Korbak et al. (2022), severely hinders the continual improvement of LLMs Zhou et al. (2024). To mitigate catastrophic forgetting, it is crucial to understand its underlying mechanisms.
Recent studies have attempted to analyze catastrophic forgetting from a “similarity” perspective Bai et al. (2024); Wu et al. (2024). Specifically, they measure the similarity between injected knowledge and previously mastered knowledge using embedding representations Mitchell et al. (2022) and knowledge graphs Tan et al. (2025); Zhu et al. (2025). Results show that knowledge highly similar to injected knowledge is more likely to be forgotten Luo et al. (2025); Li et al. (2024). While these studies provide valuable insights into catastrophic forgetting, they fail to rigorously and theoretically reveal its underlying mechanisms.
Additionally, numerous approaches have been proposed to mitigate catastrophic forgetting, including replay based Huang et al. (2024), parameter-efficient fine-tuning based Li et al. (2025), and knowledge editing based approaches Ge et al. (2024). While they alleviate forgetting to some extent, they lack a rigorous theoretical foundation and therefore cannot avoid catastrophic forgetting at theoretically.
Our work explores the mechanisms of catastrophic forgetting through both theoretical analysis and experimental validation. More specifically, we seek to answer the following three key questions regarding catastrophic forgetting:
$\bullet$ What conditions lead to catastrophic forgetting? In Section 3, we demonstrate through rigorous theoretical derivation and extensive empirical results that strongly negativee gradient similarity between injected knowledge and mastered knowledge underlies catastrophic forgetting.
$\bullet$ Which neurons are responsible for catastrophic forgetting? In Section 4, we apply gradient similarity to reveal two distinct types of neurons: conflicting neurons and collaborative neurons. Conflicting neurons induce forgetting and comprise 50%–75% of all neurons, whereas collaborative neurons alleviate forgetting and constitute 25%–50%.
$\bullet$ How to prevent catastrophic forgetting? In Section 5, we propose a novel knowledge injection method, CNL, which freezes conflicting neurons and trains collaborative neurons. CNL theoretically prevent forgetting under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set. Experiments on five models, four datasets, and four optimizers demonstrate CNL effectively mitigates forgetting under both in-set and out-of-set conditions.
To ensure the reproducibility of this study, all of the codes and datasets are included in Supplementary Material. Code and datasets are available at: https://github.com/yangmutian/Collaborative-Neuron-Learning
Figure 1: Conceptual illustration of our framework. We reveal the mechanism of catastrophic forgetting and propose a novel method to prevent its occurrence from a perspective of gradient similarity.
2 Definition of Learning and Forgetting from the Perspective of Gradient
2.1 Learning
The neural network of LLMs is defined as $f_{\theta}:\mathcal{X}\to\mathcal{Y}$, where the mapping from the input space to the output space is determined by parameters $\theta=(\theta_{1},\dots,\theta_{|\theta|})\in\mathbb{R}^{|\theta|}$.
Let $\mathcal{I}$ denote a set of knowledge samples to be injected
$\mathcal{I}=\left\{(x_{1}^{\mathcal{I}},y_{1}^{\mathcal{I}}),\dots,(x_{|\mathcal{I}|}^{\mathcal{I}},y_{|\mathcal{I}|}^{\mathcal{I}})\right\}$ (1)
Loss on a single injection sample $(x,y)\in\mathcal{I}$ is expressed as $L^{(x,y)}(\theta)=\ell(f_{\theta}(x),y)$, where $\ell(\cdot,\cdot)$ denotes the loss function. The total loss on Injection Set $\mathcal{I}$ is expressed as the mean loss over all samples
$\mathcal{L}^{\mathcal{I}}(\theta)$ $=\frac{1}{|\mathcal{I}|}\sum_{(x,y)\in\mathcal{I}}L^{(x,y)}(\theta)$
$=\frac{1}{|\mathcal{I}|}\sum_{(x,y)\in\mathcal{I}}\ell(f_{\theta}(x),y)$ (2)
Let $\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\in\mathbb{R}^{|\theta|\times 1}$ denote the gradient of $\mathcal{L}^{\mathcal{I}}(\theta)$ with respect to $\theta$. Under gradient descent with a learning rate $\eta$, the parameter update is given by
$\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (3)
Assuming an infinitesimal learning rate $\eta$, the loss after parameter update can be approximated using a first-order Taylor expansion
$\mathcal{L}^{\mathcal{I}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
$=\frac{1}{|\mathcal{I}|}\sum_{(x,y)\in\mathcal{I}}[L^{(x,y)}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)]$
$\approx\frac{1}{|\mathcal{I}|}\sum_{(x,y)\in\mathcal{I}}\left[L^{(x,y)}(\theta)-\eta\nabla_{\theta}L^{(x,y)}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right]$
$=\mathcal{L}^{\mathcal{I}}(\theta)-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (4)
Therefore, the change in loss is:
$\Delta\mathcal{L}^{\mathcal{I}}$ $=\mathcal{L}^{\mathcal{I}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)-\mathcal{L}^{\mathcal{I}}(\theta)$
$=-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (5)
Since $\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ is column vector, $\nabla_{\theta}\mathcal{I}^{\mathcal{I}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ represents the inner product of the gradient with itself. We define the Gradient Similarity as
$S(\mathcal{I},\mathcal{I})=\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (6)
Due to the non-negativity of the inner product, we have $\Delta\mathcal{L}^{\mathcal{I}}=-\eta S(\mathcal{I},\mathcal{I})\leq 0$. In practice, parameter updates are rarely zero strictly, leading to
$\Delta\mathcal{L}^{\mathcal{I}}=-\eta S(\mathcal{I},\mathcal{I})<0$ (7)
This derivation defines the process of learning: under a sufficiently small learning rate $\eta$, knowledge injection monotonically decreases the loss on Injection Set $\mathcal{I}$.
2.2 Forgetting
Subsequently, we analyze the impact of knowledge injection on samples with mastered knowledge, denoted as Mastered Set
$\mathcal{M}=\left\{(x_{1}^{\mathcal{M}},y_{1}^{\mathcal{M}}),\dots,(x_{|\mathcal{M}|}^{\mathcal{M}},y_{|\mathcal{M}|}^{\mathcal{M}})\right\}$ (8)
The loss on Mastered Set is defined as $\mathcal{L}^{\mathcal{M}}(\theta)$. After parameter update according to Eq. (3), the loss becomes $\mathcal{L}^{\mathcal{M}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$. Assuming an infinitesimal learning rate $\eta$, the expression is expanded through first-order Taylor approximation:
$\mathcal{L}^{\mathcal{M}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
$=\frac{1}{|\mathcal{M}|}\sum_{(x,y)\in\mathcal{M}}L^{(x,y)}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
$\approx\frac{1}{|\mathcal{M}|}\sum_{(x,y)\in\mathcal{M}}\left[L^{(x,y)}(\theta)-\eta\nabla_{\theta}L^{(x,y)}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right]$
$=\mathcal{L}^{\mathcal{M}}(\theta)-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (9)
The change in loss is expressed as
$\Delta\mathcal{L}^{\mathcal{M}}$ $=\mathcal{L}^{\mathcal{M}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)-\mathcal{L}^{\mathcal{M}}(\theta)$
$=-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (10)
We define the Gradient Similarity between Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ as
$S(\mathcal{M},\mathcal{I})=\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (11)
Therefore, during knowledge injection, the loss change on Mastered Set is given by
$\Delta\mathcal{L}^{\mathcal{M}}=-\eta S(\mathcal{M},\mathcal{I})$ (12)
Eq. (12) reveals a critical insight: the loss on the Mastered Set increases when the gradient similarity is negative. Therefore, mastered samples with more negative gradient similarity (i.e., stronger alignment in the opposite direction) are more susceptible to catastrophic forgetting.
3 Gradient Similarity Measures Catastrophic Forgetting
To validate the conclusion of Eq. (12), experiments on five models and four datasets were conducted.
3.1 Experiment Procedure
Datasets: We employed standard benchmarks including Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020), Medical Question Answering (MedQA) Jin et al. (2021), AI2 Reasoning Challenge (ARC-C) Clark et al. (2018), and Commonsense Question Answering (CSQA) Talmor et al. (2019). The detailed information of datasets is shown in Appendix A.
Models: The experiments were conducted on Qwen Yang et al. (2024) and LLaMA Grattafiori et al. (2024) families, including Qwen2.5 1.5B, Qwen2.5 3B, Qwen2.5 7B, LLaMA3.2 1B, and LLaMA3.2 3B.
3.2 Catastrophic Forgetting Happens during Knowledge Injection
To distinguish Injection Set $\mathcal{I}$ and Mastered Set $\mathcal{M}$, LLMs were asked to answer the questions in datasets (inference details in Appendix B). Correctly answered questions were categorized into $\mathcal{M}$, while incorrect ones formed $\mathcal{I}$. Appendix C exhibits the distribution of Injection Set $\mathcal{I}$ and Mastered Set $\mathcal{M}$.
Subsequently, LLMs were trained on Injection Set $\mathcal{I}$ for knowledge injection (training details in Appendix  D). The correctly answered samples in Injection Set $\mathcal{I}$ were used to evaluate learning, while the incorrectly answered samples in Mastered Set $\mathcal{M}$ were used to measure forgetting. Results in Appendix E reveal that while LLMs successfully learn new knowledge from Injection Set $\mathcal{I}$, they simultaneously suffer forgetting on Mastered Set $\mathcal{M}$.
Notably, we extended our evaluation to parameter-efficient fine-tuning (PEFT) methods and observed that techniques like Low-Rank Adaptation (LoRA) Hu et al. (2022) could not mitigate this catastrophic forgetting. Detailed experimental settings and results are provided in Appendix F.
3.3 Strongly Negative Gradient Similarity Leads to Forgetting
Ultimately, gradient similarities between samples in Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ were computed. In light of Eq. (12), samples with negative gradient similarity were ranked by the magnitude of their similarity. The top $1/3$ with the largest magnitudes were assigned to the “Sim” group, representing strongly negative similarity, while the bottom $1/3$ were assigned to the “Dissim” group.
Results in Table 1 indicate that catastrophic forgetting is predominantly concentrated in the “Sim” group. This empirical evidence provides strong support for Eq. (12), confirming that strongly negative gradient similarity drives catastrophic forgetting.
Table 1: Relationship between gradient similarity and catastrophic forgetting. Results are reported in the format of “Number of forgotten questions (Percentage)”. Samples with negative gradient similarity were ranked by magnitude: the top $1/3$ with the largest magnitudes were labeled “Sim” (High Negative Similarity), while the bottom $1/3$ were labeled “Dissim”. Results indicate that catastrophic forgetting occurs primarily in the “Sim” group.
Model MMLU MedQA ARC-C CSQA
Dissim Sim Dissim Sim Dissim Sim Dissim Sim
Qwen 1.5B 4 137 37 113 3 72 10 102
(1.4%) (48.9%) (20.6%) (62.8%) (1.0%) (23.2%) (3.7%) (38.1%)
Qwen 3B 0 118 0 113 7 193 0 84
(0.0%) (37.9%) (0.0%) (54.6%) (2.0%) (56.4%) (0.0%) (31.5%)
Qwen 7B 3 120 3 116 0 17 0 40
(0.9%) (36.0%) (1.2%) (46.0%) (0.0%) (5.1%) (0.0%) (13.7%)
LLaMA 1B 0 68 0 43 6 91 9 89
(0.0%) (58.6%) (0.0%) (42.2%) (6.6%) (100%) (10.1%) (100.0%)
LLaMA 3B 0 83 28 32 1 73 0 34
(0.0%) (33.2%) (11.1%) (12.6%) (0.4%) (25.9%) (0.0%) (13.6%)
4 Conflicting Neurons Cause Catastrophic Forgetting
While LLMs exhibit overall forgetting on Mastered Set, it remains unclear whether forgetting is driven by all neurons or only a subset. Herein, gradient similarity is employed to identify the neurons contributing to forgetting. It should be clarified that in this paper, “neurons” and “parameters” carry essentially the same meaning. We use parameters when emphasizing their vector form for computation, and neurons when emphasizing their count.
The gradient similarity of the $j$-th neuron $\theta_{j}$ between Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ is defined as
$s_{\theta_{j}}(\mathcal{M},\mathcal{I})=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)$ (13)
$\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)$ and $\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)$ represent the components of gradients on $\theta_{j}$ for Mastered and Injection Set, respectively.
The global gradient similarity of LLMs is equivalent to the sum of gradient similarities of individual neurons. Therefore, Eq. (11) is expressed as
$S(\mathcal{M},\mathcal{I})$ $=\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$
$=\sum_{j=1}^{|\theta|}\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
$=\sum_{j=1}^{|\theta|}s_{\theta_{j}}(\mathcal{M},\mathcal{I})$ (14)
According to Eq. (12), if the gradient similarity of neuron $\theta_{j}$ with respect to Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ is negative, this neuron contributes to an increase in loss and induce forgetting. We refer to such neurons as Conflicting Neurons. The set of conflicting neurons is defined as
$\theta_{CF}:=\left\{\theta_{j}\mid s_{\theta_{j}}(\mathcal{M},\mathcal{I})<0\right\}$ (15)
Conversely, if the gradient similarity of neuron $\theta_{j}$ is positive, this neuron contributes to a decrease in loss and mitigate forgetting. Such neurons are refered as Collaborative Neurons. The set of collaborative neurons is defined as
$\theta_{CB}:=\left\{\theta_{j}\mid s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0\right\}$ (16)
Table 2 presents that nearly 25%-50% of neurons contribute to positive gradient similarity and are termed collaborative neurons, while 50%-75% of neurons contribute to negative gradient similarity and are termed conflicting neurons. These collaborative neurons are insufficient to counteract the effect of conflicting neurons, thereby leading to the phenomenon of catastrophic forgetting in LLMs.
This result reveals the coexistence of two types of neurons during catastrophic forgetting, which provide theoretical foundation for the following knowledge injection method.
Table 2: Distribution of collaborative neurons (COLL) and conflicting neurons (CONF). "PROP" represents the proportion of a specific neuron type relative to the total number of neurons. “GRAD” represents the gradient similarity sum for that neuron type. “TOTAL” represents the gradient similarity sum for all neurons. Typically, collaborative neurons account for 25%–50% of all neurons, while conflicting neurons account for 50%–75%.
Model Metric MMLU MedQA ARC-C CSQA
Coll Conf Coll Conf Coll Conf Coll Conf
Qwen 1.5B Prop 47.7% 52.3% 42.1% 57.9% 49.7% 50.3% 48.4% 51.6%
Grad +692 -1851 +567 -3186 +635 -1021 +577 -1438
Total -1159 -2619 -386 -861
Qwen 3B Prop 49.4% 50.6% 44.6% 55.4% 51.7% 48.3% 52.9% 47.1%
Grad +1626 -2979 +1538 -4319 +2237 -2569 +2703 -3276
Total -1353 -2781 -332 -573
Qwen 7B Prop 49.6% 50.4% 49.0% 51.0% 51.2% 48.8% 51.6% 48.4%
Grad +710 -1242 +661 -1282 +433 -664 +687 -940
Total -532 -621 -231 -253
LLaMA 1B Prop 34.1% 65.9% 24.0% 76.0% 28.6% 71.4% 42.1% 57.9%
Grad +2116 -64263 +54 -116388 +373 -88549 +4349 -41111
Total -62147 -116334 -88176 -36762
LLaMA 3B Prop 37.8% 62.2% 30.4% 69.6% 46.5% 53.5% 42.7% 57.3%
Grad +282 -2528 +116 -2784 +734 -2161 +404 -2178
Total -2246 -2668 -1427 -1774
5 Preventing Catastrophic Forgetting by Freezing Conflicting Neurons
Since conflicting neurons are responsible for catastrophic forgetting, a natural question arises: can catastrophic forgetting be avoided by freezing conflicting neurons and training only collaborative neurons? We refer to this learning paradigm as Collaborative Neuron Learning (CNL).
5.1 Establishment of CNL
Table 3: Effectiveness of knowledge injection using FT and CNL. Results are reported in the format of “Number of questions (Percentage)”. FT leads to the forgetting of previously mastered knowledge comparable in magnitude to the newly injected knowledge, whereas CNL induces negligible catastrophic forgetting.
Model Method MMLU MedQA ARC-C CSQA
Learned Forgot Learned Forgot Learned Forgot Learned Forgot
Qwen 1.5B FT 168 188 191 264 104 88 86 163
(24.8%) (21.4%) (27.1%) (46.5%) (32.1%) (9.3%) (31.6%) (19.8%)
CNL 168 0 193 0 163 0 147 0
(24.8%) (0.0%) (27.4%) (0.0%) (50.3%) (0.0%) (54.0%) (0.0%)
Qwen 3B FT 114 126 119 160 71 286 77 105
(19.5%) (13.0%) (18.6%) (25.3%) (33.5%) (27.0%) (31.7%) (12.3%)
CNL 116 0 119 0 76 0 81 0
(19.8%) (0.0%) (18.6%) (0.0%) (35.8%) (0.0%) (33.3%) (0.0%)
Qwen 7B FT 127 193 110 169 62 19 68 48
(28.0%) (17.5%) (21.9%) (21.9%) (46.3%) (1.7%) (34.5%) (5.3%)
CNL 189 0 249 0 90 0 144 0
(41.6%) (0.0%) (49.6%) (0.0%) (67.2%) (0.0%) (73.1%) (0.0%)
LLaMA 1B FT 158 73 105 43 328 154 333 176
(14.0%) (17.0%) (11.0%) (13.4%) (34.4%) (48.7%) (47.8%) (44.1%)
CNL 201 0 112 0 335 0 338 0
(17.9%) (0.0%) (11.8%) (0.0%) (35.1%) (0.0%) (48.5%) (0.0%)
LLaMA 3B FT 133 96 113 101 85 76 60 35
(18.3%) (11.6%) (22.2%) (13.2%) (24.5%) (8.2%) (19.5%) (4.4%)
CNL 133 0 125 0 86 0 63 0
(18.3%) (0.0%) (24.6%) (0.0%) (24.8%) (0.0%) (20.5%) (0.0%)
To indicate the sign of gradient similarity for the $j$-th neuron, we define an indicator function $I(\cdot)$ as
$I\!\left(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0\right)=\begin{cases}1,&\text{if }s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0\\ 0,&\text{if }s_{\theta_{j}}(\mathcal{M},\mathcal{I})<0\end{cases}$ (17)
Applying the indicator function to all neurons yields
$\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]$ (18)
$=\big(I(s_{\theta_{1}}(\mathcal{M},\mathcal{I})\geq 0),\ldots,I(s_{\theta_{|\theta|}}(\mathcal{M},\mathcal{I})\geq 0)\big)^{\top}$
CNL freezes neurons with negative gradient similarity, leading to a parameter update rule defined as
$\theta\leftarrow\theta-\eta\,\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]\odot\nabla_{\theta}L^{\mathcal{I}}(\theta)$ (19)
where $\odot$ denotes the Hadamard (element-wise) product. According to Eq. (2.2), the change in loss on Mastered Set $\mathcal{M}$ after parameter update is given by
$\Delta L^{\mathcal{M}}(\theta)$
$=-\eta\,\nabla_{\theta}L^{\mathcal{M}}(\theta)^{\top}\big(\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]\odot\nabla_{\theta}L^{\mathcal{I}}(\theta)\big)$
$=-\eta\sum_{j=1}^{|\theta|}\Big[\nabla_{\theta_{j}}L^{\mathcal{M}}(\theta)\,I(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0)\,\nabla_{\theta_{j}}L^{\mathcal{I}}(\theta)\Big]$
$=-\eta\sum_{j=1}^{|\theta|}\Big[I(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0)\,s_{\theta_{j}}(\mathcal{M},\mathcal{I})\Big]$ (20)
In practical settings, it is almost impossible for all parameter gradients to be exactly zero. Therefore, we have
$\Delta L^{\mathcal{M}}(\theta)$ (21)
$=-\eta\sum_{j=1}^{|\theta|}\Big[I(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0)\,s_{\theta_{j}}(\mathcal{M},\mathcal{I})\Big]\;<0$
This result indicates that, under the assumption of an infinitesimally small learning rate $\eta$, CNL guarantees a monotonic decrease of the loss on Mastered Set $\mathcal{M}$. In other words, CNL theoretically prevents catastrophic forgetting.
To experimentally evaluate the effectiveness of CNL, we performed knowledge injection using both CNL and conventional fine tuning (FT). Their forgetting was compared under the constraint that CNL learned at least as many new samples as FT. As shown in Table 3, FT forgets a number of mastered samples comparable to the amount of newly injected knowledge, whereas CNL achieve zero forgetting.
Although CNL exhibits excellent performance in preventing forgetting, we have to point out that CNL achieves zero forgetting only under two restrictive conditions: (i) the learning rate $\eta$ is sufficiently small, and (ii) Mastered Set $\mathcal{M}$ is explicitly known. The case where Mastered Set $\mathcal{M}$ is not explicitly available will be discussed in Section 5.4.
5.2 Extending CNL to More Complex Optimizers
Table 4: Comparison of knowledge injection performance across different optimizers. Qwen2.5 1.5B is selected for experiment. Whether using SGD, Momentum, Adam, or AdamW, CNL consistently achieves zero catastrophic forgetting while maintaining effective knowledge injection.
Optimizer Method MMLU MedQA ARC-C CSQA
Learned Forgot Learned Forgot Learned Forgot Learned Forgot
SGD FT 168 188 191 264 104 88 86 163
(24.8%) (21.4%) (27.1%) (46.5%) (32.1%) (9.3%) (31.6%) (19.8%)
CNL 168 0 193 0 163 0 147 0
(24.8%) (0.0%) (27.1%) (0.0%) (50.3%) (0.0%) (54.0%) (0.0%)
Momentum FT 341 450 257 398 261 169 205 233
(50.3%) (51.3%) (36.5%) (70.1%) (80.6%) (17.9%) (75.4%) (28.3%)
CNL 408 0 461 0 291 0 240 0
(60.2%) (0.0%) (65.5%) (0.0%) (89.8%) (0.0%) (88.2%) (0.0%)
Adam FT 647 203 704 181 320 143 266 168
(95.4%) (23.1%) (100.0%) (31.9%) (98.8%) (15.1%) (97.8%) (20.4%)
CNL 648 0 704 0 323 0 266 0
(95.6%) (0.0%) (100.0%) (0.0%) (99.7%) (0.0%) (97.8%) (0.0%)
AdamW FT 638 195 703 174 320 143 251 294
(94.1%) (22.2%) (99.9%) (30.6%) (98.8%) (15.1%) (92.3%) (35.7%)
CNL 642 0 703 0 323 0 266 0
(94.7%) (0.0%) (99.9%) (0.0%) (99.7%) (0.0%) (97.8%) (0.0%)
After establishing CNL with Stochastic Gradient Descent (SGD) optimizer Rumelhart et al. (1985), we further extend CNL to Momentum Rumelhart et al. (1985), Adam Kingma (2014), AdamW Loshchilov and Hutter (2019), and arbitrary optimizer.
In Appendix G, rigorous proofs are provided that the loss on Mastered Set decreases monotonically, as long as the parameter update of Momentum, Adam, AdamW, arbitrary optimizer follows Eq. (27), (38), (43), and (58).
Table 4 demonstrates the experimental results of CNL with difference optimizers. When FT is applied, all four optimizers suffer from severe catastrophic forgetting. In contrast, CNL effectively prevents catastrophic forgetting across all four optimizers. In particular, when using Adam optimizer, CNL is able to successfully inject nearly all (95.6%-100.0%) required knowledge while exhibiting zero forgetting.
Additionally, the dynamic characteristic of learning and forgetting is analyzed in Appendix. H. Although CNL freezes conflicting neurons and trains only collaborative neurons, its learning efficiency remains comparable to that of conventional FT. This further demonstrates the engineering potential of our CNL.
5.3 Extending CNL to More Complex Datasets
To further valid CNL on more complex datasets, MMLU, MedQA, ARC-C, and CSQA were merged into a unified dataset, referred to as MMAC. The performance of FT and CNL on MMAC is studied in Appendix I.
Results demonstrate that, while successfully injecting 95% knowledge into LLMs, FT leads to 91.7% forgetting of mastered knowledge, while CNL exhibits zero forgetting. These findings further demonstrate that CNL can effectively prevent catastrophic forgetting even under more complex and large-scale data conditions.
5.4 Extending CNL to Out-of-Set Generalization
We have verified that CNL achieves zero forgetting under in-set setting. However, it is hard to explicitly obtain all Mastered Set and Injection Set in practice. Therefore, the out-of-set generalization capability of CNL is analyzed.
First, Injection Set $\mathcal{I}$ was split into two disjoint subsets in a $1:1$ ratio, one used for training and the other reserved for evaluation. The “INJECTION" row of Table 5 demonstrates that CNL still achieves zero forgetting because the loss on Mastered Set can be explicitly calculated.
Subsequently, the same procedure is applied to the Mastered Set $\mathcal{M}$, as reported in the “MASTERED” row of Table 5. Under this setting, the loss on the evaluation Mastered subset cannot be directly computed and is instead estimated using the training Mastered subset. Due to this estimation error, CNL cannot guarantee a monotonic decrease in the evaluation loss. Nevertheless, while zero forgetting cannot be fully achieved, CNL still substantially alleviates catastrophic forgetting, reducing it by 59.1%–81.7% compared to conventional FT.
Furthermore, the training Mastered subset was combined with Injection Set during training to mitigate forgetting, a strategy known as replay. RP follows the same training setup as FT. Results in the “RP" row of Table 5 demonstrate that CNL not only learns more new knowledge but also forgets less mastered knowledge compared to replay.
These results demonstrate that CNL not only achieves zero forgetting under the ideal in-set condition, but also significantly alleviates catastrophic forgetting under more practical out-of-set evaluation settings.
6 Related Work
Table 5: Performance evaluation under out-of-set setting. Qwen2.5 1.5B is selected for experiment. To assess generalization, both Injection Set and Mastered Set were partitioned into training and evaluation subsets. This table reports the learning and forgetting performance on unseen evaluation data, demonstrating the excellent generalization capability of CNL.
Method MMLU MedQA ARC-C CSQA
Learned Forgot Learned Forgot Learned Forgot Learned Forgot
Injection FT 24 48 34 75 29 71 22 114
(7.3%) (5.5%) (9.7%) (13.2%) (17.9%) (7.5%) (16.2%) (13.8%)
CNL 31 0 39 0 29 0 23 0
(9.5%) (0.0%) (11.1%) (0.0%) (17.9%) (0.0%) (16.9%) (0.0%)
Mastered FT 178 110 208 143 104 49 86 94
(26.3%) (26.0%) (29.5%) (50.4%) (32.1%) (10.4%) (31.6%) (22.8%)
RP 157 61 124 58 86 14 70 28
(23.2%) (14.4%) (17.6%) (20.4%) (26.5%) (3.0%) (25.7%) (6.8%)
CNL 179 46 243 33 116 9 147 18
(26.4%) (10.9%) (34.5%) (11.6%) (35.8%) (1.9%) (54.0%) (4.4%)
6.1 Knowledge Injection
Existing knowledge injection methods are broadly categorized into two types: parametric methods and non-parametric methods Doan et al. (2021); Kotha et al. (2023).
Parametric methods directly modify model parameters to encode target knowledge via instruction tuning or continued pre-training, but often suffer from catastrophic forgetting of previously acquired knowledge Singhal et al. (2023); Han et al. (2023); Qiu et al. (2024). Non-parametric methods, such as retrieval-augmented generation, augment model inference with externally retrieved information without directly modifying model parameters Lewis et al. (2020); Ram et al. (2023); Xiong et al. (2024). As a result, they mitigate parameter-level interference, but their effectiveness is constrained by retrieval quality and limited access to implicit knowledge required for complex reasoning Salemi and Zamani (2024); Zhu et al. (2024); Kim et al. (2025) .
6.2 Mitigating Catastrophic Forgetting
Approaches to mitigate catastrophic forgetting have been extensively studied De Lange et al. (2021). Existing strategies generally operate along three dimensions.
Data-based approaches, such as experience replay, mix mastered knowledge with injected knowledge to preserve historical gradient signals, but suffer from storage overhead and privacy concerns. Constraint-based approaches, including PEFT, reduce forgetting by restricting the scope of parameter updates or penalizing large deviations from pretrained weights Wang et al. (2021); Hu et al. (2022). Localization-based approaches, such as knowledge editing methods, attempt to identify and modify task-relevant representation regions to achieve targeted knowledge updates with minimal side effects Meng et al. (2022, 20222).
Although effective in practice, most existing methods rely on heuristic designs and lack an explicit characterization of when and why forgetting occurs. In contrast, this work seeks to establish a principled connection between forgetting behavior and optimization dynamics.
6.3 Gradient Analysis in Continual Learning
From an optimization perspective, the gradient inner product (or gradient similarity) serves as a fundamental indicator of interference between learning objectives Yu et al. (2020); Liu et al. (2021). When gradients induced by new knowledge conflict with those associated with previously learned knowledge, optimization inevitably degrades prior performance, manifesting as catastrophic forgetting Farajtabar et al. (2020). Previous studies in continual and lifelong learning have leveraged gradient projection or alignment techniques to mitigate such conflicts at a global level Wang et al. (2025).
This work advances this perspective by decomposing gradient inner products at the neuron level, enabling identification of conflicting and collaborative components within the model. By selectively suppressing neuron-level conflicting gradients while preserving collaborative ones, forgetting mitigation is transformed from heuristic data balancing into a theoretically grounded training rule that is compatible with standard optimizers.
Conclusion
Catastrophic forgetting during knowledge injection poses a fundamental challenge to the continual learning capability of LLMs. In this work, we establish a gradient-based theoretical framework to explain the underlying mechanism of catastrophic forgetting. We show that strongly negative gradient similarity is the fundamental cause of forgetting and further leverage this insight to categorize neurons into two types: conflicting neurons, which induce forgetting and account for 50%–75% of all neurons, and collaborative neurons, which mitigate forgetting and account for the remaining 25%–50%.
Based on this analysis, we propose a novel knowledge injection method, CNL, which freezes conflicting neurons and updates only collaborative neurons. We theoretically show that, under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set, CNL prevents catastrophic forgetting. Extensive experiments on five LLMs, four datasets, and four optimizers further validate our analysis. CNL achieves zero forgetting under in-set evaluation and reduces catastrophic forgetting by 59.1%–81.7% under more practical out-of-set settings.
Overall, this work provides both theoretical insights and practical solutions for mitigating catastrophic forgetting in knowledge injection, offering a principled foundation for continual knowledge acquisition in LLMs.
Acknowledgments
This work was supported by the Noncommunicable Chronic Diseases-National Science and Technology Major Project under Grants 2024ZD0522702 and 2024ZD0528306, and the National Natural Science Foundation of China under Grant 62433007.
Appendix A Dataset Details
To verify our conclusions, we selected four datasets for experimentation: MMLU, MedQA, ARC-C, and CSQA. The data processing method is described as follows:
  • MMLU: 1/10 subsampling is performed to result in 1, 555 remaining questions.
  • MedQA: 1/10 subsampling is performed to result in 1, 273 remaining questions.
  • ARC-C: 1/2 subsampling is performed to result in 1, 270 remaining questions.
  • CSQA: 1/10 subsampling is performed to result in 1, 096 remaining questions.
Appendix B Model Inference Details
Model inference was conducted in a Python (3.12.2) environment, with the main packages including transformers (4.48.0) and torch (2.2.0). To ensure experimental reproducibility, the temperature was set to 0. Model inference precision was set to FP32, and experiments were conducted on a single A800 GPU.
Appendix C Model Inference Results
To identify Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$, LLMs were asked to answer the questions in the datasets. Correctly answered questions were assigned to Mastered Set $\mathcal{M}$, while incorrectly answered questions were assigned to Injection Set $\mathcal{I}$. The distribution of different sets are described in Table. 6.
Table 6: Number of questions in Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$. “Injection” denotes the number of samples in Injection Set (answered incorrectly), while “Mastered” denotes the number of samples in Mastered Set (answered correctly).
Model MMLU MedQA ARC-C CSQA
Injection Mastered Injection Mastered Injection Mastered Injection Mastered
Qwen 1.5B 678 877 704 568 324 946 272 824
Qwen 3B 585 970 639 633 212 1058 243 853
Qwen 7B 454 1101 502 770 134 1136 197 899
LLaMA 1B 1126 429 952 320 954 316 697 399
LLaMA 3B 727 828 508 764 347 923 308 788
Appendix D Model Training Details
The environment of model training was essentially the same as that for model inference. The number of epochs was set to 25. For the SGD optimizer, the learning rate included three values according to their magnitude of gradient: $2\text{e-}9$, $2\text{e-}7$, and $1\text{e-}7$ (see the code for exact configurations). For Momentum optimizer, the learning rate was consistent with SGD, and the momentum coefficient was set to 0.9. For Adam optimizer (without regularization), $\beta_{1}$ and $\beta_{2}$ were 0.9 and 0.999 respectively, and $\epsilon$ was $1\text{e-}8$. For AdamW optimizer, the regularization parameter (weight decay) was $1\text{e-}1$. Model training was performed in FP32 precision on a single NVIDIA A800 GPU.
Appendix E Model Training Results
The samples in Injection Set $\mathcal{I}$ were trained for knowledge injection. After training, the number of correctly answered questions in Injection Set $\mathcal{I}$ was employed to evaluate learning, while the number of incorrectly answered questions in Mastered Set $\mathcal{M}$ was employed to evaluate forgetting. The results in Table. 7 indicates that serious catastrophic forgetting happens during knowledge injection.
Table 7: Learning and forgetting during knowledge injection. Results are presented in the format of “Number of Items (Percentage)”. After training, although new knowledge is successfully injected into LLMs, severe catastrophic forgetting is also observed.
Model MMLU MedQA ARC-C CSQA
Learned Forgot Learned Forgot Learned Forgot Learned Forgot
Qwen 1.5B 168 188 191 264 104 88 86 163
(24.8%) (21.4%) (27.1%) (46.5%) (32.1%) (9.3%) (31.6%) (19.8%)
Qwen 3B 114 126 119 160 71 286 77 105
(19.5%) (13.0%) (18.6%) (25.3%) (33.5%) (27.0%) (31.7%) (12.3%)
Qwen 7B 127 193 110 169 62 19 68 48
(28.0%) (17.5%) (21.9%) (21.9%) (46.3%) (1.7%) (34.5%) (5.3%)
LLaMA 1B 158 73 105 43 328 154 333 176
(14.0%) (17.0%) (11.0%) (13.4%) (34.4%) (48.7%) (47.8%) (44.1%)
LLaMA 3B 133 96 113 101 85 76 60 35
(18.3%) (11.6%) (22.2%) (13.2%) (24.5%) (8.2%) (19.5%) (4.4%)
Appendix F Evaluation on Parameter-Efficient Fine-Tuning
We evaluate LoRA as a representative parameter-efficient fine-tuning (PEFT) baseline.
All experiments are conducted on the Qwen2.5-1.5B model. For a fair comparison with standard fine-tuning (Appendix D), we adopt the same training protocol, using the SGD optimizer with a learning rate of $\eta=1\times 10^{-5}$. The LoRA configuration employs a rank of $r=16$ and a scaling factor of $\alpha=32$, with LoRA modules applied to all linear layers in the attention mechanism. The batch size and number of training epochs are identical to those used in Appendix D.
As shown in Table 8, LoRA achieves learning performance comparable to full fine-tuning, but does not alleviate catastrophic forgetting, exhibiting similarly severe forgetting across all evaluated datasets.
Table 8: Comparison of FT and LoRA under during knowledge injection. Forgetting cannot be effectively prevented when LoRA is employed.
Method MMLU MedQA ARC-C CSQA
Learned Forgot Learned Forgot Learned Forgot Learned Forgot
FT 168 188 191 264 104 88 86 163
(24.8%) (21.4%) (27.1%) (46.5%) (32.1%) (9.3%) (31.6%) (19.8%)
LoRA 168 176 193 286 104 79 86 164
(24.8%) (20.1%) (27.4%) (50.4%) (32.1%) (8.4%) (31.6%) (19.9%)
Appendix G Theoretical Derivation for More Complex Optimizers
G.1 Momentum Optimizer
When performing gradient descent with learning rate $\eta$, the update rule for parameters $\theta$ in the Momentum optimizer is as follows:
$m$ $\leftarrow\beta m+\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (22)
$\theta$ $\leftarrow\theta-\eta m$
where $m$ performs an exponential weighted average of historical gradients for accelerated convergence. $\beta\in[0,1)$ is the momentum coefficient, used to control the decay rate of historical momentum information.
The impact of parameter updates on existing knowledge is analyzed. Similar to Eq. (2.2), we perform a Taylor expansion on loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over Mastered Set
$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta m)$ (23)
$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta m)$
The change in loss on Mastered Set is:
$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}m}_{S^{mom}(\mathcal{M},\mathcal{I})}$ (24)
Here, $S_{mom}(\mathcal{M},\mathcal{I})$ is defined as the momentum similarity. If $S_{mom}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting. Utilizing the additivity of inner product, we decompose the global momentum similarity onto each neuron
$S^{mom}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{mom}$ (25)
where $s_{\theta_{j}}^{mom}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot m_{j}$. Based on the sign of $s_{\theta_{j}}^{mom}$, we classify the neurons into two categories:
$\theta_{CF}$ $=\left\{\theta_{j}\mid s_{\theta_{j}}^{mom}(\mathcal{M},\mathcal{I})<0\right\}$ (Conflicting Neurons) (26)
$\theta_{CO}$ $=\left\{\theta_{j}\mid s_{\theta_{j}}^{mom}(\mathcal{M},\mathcal{I})\geq 0\right\}$ (Collaborative Neurons)
Similar to CNL with SGD, the parameter update rule of CNL for momentum is
$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{mom}\geq 0)\odot m\right]$ (27)
Under this rule, the loss in Master Set is
$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{mom}\geq 0)\odot m\right]$ (28)
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{mom}\geq 0)\cdot(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot m_{j})\right]$
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{mom}\geq 0)\cdot s_{\theta_{j}}^{mom}\right]\leq 0$
In practical settings, it is nearly impossible for all gradients to be exactly zero. Therefore we get
$\Delta\mathcal{L}^{\mathcal{M}}<0$ (29)
G.2 Adam Optimizer (without weight decay)
The conventional update rule for the Adam optimizer is as follows:
$m$ $\leftarrow\beta_{1}m+(1-\beta_{1})g_{t}$ (30)
$v$ $\leftarrow\beta_{2}v+(1-\beta_{2})g_{t}^{2}\quad$
$\hat{m}$ $=\frac{m}{1-\beta_{1}^{t}},\quad\hat{v}=\frac{v}{1-\beta_{2}^{t}}\quad(\text{bias correction})$
$\theta$ $\leftarrow\theta-\eta\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}$
where $g_{t}=\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ represents the gradient vector at step $t$. $m$ and $v$ are the exponential moving average estimates of the first moment (mean) and the second raw moment (uncentered variance) of the gradients, respectively. $\beta_{1},\beta_{2}\in[0,1)$ are the exponential decay rates for the first and second moment estimates, respectively. $\hat{m}$ and $\hat{v}$ are the bias-corrected moment estimates, intended to counteract the bias caused by initializing $m,v$ to zero during the initial training phase. $\epsilon$ is a numerical stability constant added to prevent division by zero.
For notational convenience, we denote the actual update direction calculated by Adam as the vector $U_{adam}$
$U^{adam}=\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}$ (31)
Therefore, the conventional update rule for the Adam optimizer is
$\theta=\theta-\eta U^{adam}.$ (32)
Next, we analyze the impact of parameter updates on existing knowledge. We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set:
$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta U^{adam})$ (33)
$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta U^{adam})$
The change in loss on the mastered set is:
$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}U^{adam}}_{S^{adam}(\mathcal{M},\mathcal{I})}$ (34)
Here, $S_{adam}(\mathcal{M},\mathcal{I})$ is defined as the Adam similarity. If $S_{adam}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting.
Utilizing the additivity of the inner product, we decompose the global Adam similarity onto each neuron:
$S^{adam}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{adam}$ (35)
where the similarity component of the $j$-th neuron is:
$s_{\theta_{j}}^{adam}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adam}_{j}$ (36)
We classify the neurons into two categories:
$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adam}<0\}\quad\text{(Conflicting Neurons)}$ (37)
$\theta_{CO}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adam}\geq 0\}\quad\text{(collaborative Neurons)}$
The modified CNL parameter update formula is:
$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{adam}\geq 0)\odot U^{adam}\right]$ (38)
Substituting the CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{adam}\geq 0)\odot U^{adam}\right]$ (39)
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{adam}\geq 0)\cdot\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adam}_{j}\right)\right]$
$=-\eta\sum_{j=1}^{|\theta|}[I(s_{\theta_{j}}^{adam}\geq 0)\cdot s_{\theta_{j}}^{adam}]\leq 0$
In practical engineering,
$\Delta\mathcal{L}^{\mathcal{M}}<0$ (40)
G.3 AdamW Optimizer
The AdamW optimizer introduces decoupled weight decay on top of Adam, separating the weight decay term from the gradient update. Its update rule is as follows:
$m$ $\leftarrow\beta_{1}m+(1-\beta_{1})g_{t}$ (41)
$v$ $\leftarrow\beta_{2}v+(1-\beta_{2})g_{t}^{2}$
$\hat{m}$ $=\frac{m}{1-\beta_{1}^{t}},\quad\hat{v}=\frac{v}{1-\beta_{2}^{t}}$
$\theta$ $\leftarrow\theta-\eta\left(\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}+\lambda\theta\right)$
where the definitions of $m,v,\hat{m},\hat{v},\beta_{1},\beta_{2},\epsilon$ are consistent with the Adam optimizer. $\lambda$ represents the Weight Decay Coefficient. Note that the update vector $U_{adamw}$ here includes not only the gradient-based adaptive step size but also the regularization term $\lambda\theta$ pointing towards the origin.
For notational convenience, we denote the actual total update direction calculated by AdamW (including weight decay) as the vector $U^{adamw}$:
$U^{adamw}=\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}+\lambda\theta$ (42)
The rule of parameter update becomes
$\theta=\theta-\eta U^{adamw}.$ (43)
Next, we analyze the impact of parameter updates on existing knowledge. We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set:
$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta U^{adamw})$ (44)
$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta U^{adamw})$
The change in loss on Mastered Set is:
$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}U^{adamw}}_{S^{adamw}(\mathcal{M},\mathcal{I})}$ (45)
Here, $S_{adamw}(\mathcal{M},\mathcal{I})$ is defined as the AdamW similarity. If $S_{adamw}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting.
Utilizing the additivity of the inner product, we decompose the global AdamW similarity onto each neuron
$S^{adamw}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{adamw}$ (46)
where the similarity component of the $j$-th neuron is:
$s_{\theta_{j}}^{adamw}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adamw}_{j}$ (47)
We classify the neurons into two categories:
$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adamw}<0\}\quad\text{(Conflicting Neurons)}$ (48)
$\theta_{CB}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adamw}\geq 0\}\quad\text{(collaborative Neurons)}$
The modified CNL parameter update formula is
$\theta\leftarrow\theta-\eta\left[I(S^{adamw}\geq 0)\odot U^{adamw}\right]$ (49)
Substituting the CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\cdot}\left[\mathbb{I}(\mathbf{S}^{adamw}\geq 0)\odot U^{adamw}\right]$ (50)
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{adamw}\geq 0)\cdot\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adamw}_{j}\right)\right]$
$=-\eta\sum_{j=1}^{|\theta|}[I(s_{\theta_{j}}^{adamw}\geq 0)\cdot s_{\theta_{j}}^{adamw}]\leq 0$
In practical engineering
$\Delta\mathcal{L}^{\mathcal{M}}<0$ (51)
G.4 General Optimizers
To derive a unified theoretical framework, we abstract the specific update rules of different optimizers into a generalized form. Regardless of whether the optimizer utilizes momentum, adaptive learning rates, or weight decay, its parameter update can ultimately be expressed as applying a specific update vector $F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ to the current parameters.
The general update rule is defined as:
$\theta$ $\leftarrow\theta-\eta F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ (52)
where $\eta$ is the global learning rate. Next, we analyze the impact of this general parameter update on existing knowledge. We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set
$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)))$ (53)
$\approx\mathcal{L}^{\mathcal{M}}(\theta)-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$
The change in loss on the mastered set is:
$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))}_{S^{opt}(\mathcal{M},\mathcal{I})}$ (54)
Utilizing the additivity of the inner product, we decompose the global similarity onto each neuron $j$:
$S^{opt}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{opt}$ (55)
where the similarity component of the $j$-th neuron is determined by the projection of the optimizer’s update scalar:
$s_{\theta_{j}}^{opt}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot F^{opt}_{j}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ (56)
Based on the sign of $s_{\theta_{j}}^{opt}$, we universally classify the neurons into two categories for any optimizer:
$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{opt}<0\}\quad\text{(Conflicting Neurons)}$ (57)
$\theta_{CB}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{opt}\geq 0\}\quad\text{(Collaborative Neurons)}$
The generalized CNL parameter update formula is:
$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{opt}\geq 0)\odot F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$ (58)
Substituting the generalized CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{opt}\geq 0)\odot F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$ (59)
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{opt}\geq 0)\cdot(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot F^{opt}_{j}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{opt}\geq 0)\cdot s_{\theta_{j}}^{opt}\right]\leq 0$
In practical engineering, we have:
$\Delta\mathcal{L}^{\mathcal{M}}<0$ (60)
Appendix H Dynamic Characteristic of CNL
The dynamic characteristics of learning and forgetting during knowledge injection are shown in Fig. 2. Although CNL freezes conflicting neurons, its training efficiency does not significantly lag behind that of conventional FT.
MMLU
MedQA
ARC-C
CSQA
SGD
Learning
Forgetting
Momentum
Learning
Forgetting
Adam
Learning
Forgetting
AdamW
Learning
Forgetting
Figure 2: Learning and forgetting curves across four datasets and four optimizers. The rows represent SGD, Momentum, Adam, and AdamW. The inner labels indicate the metric type (Learning vs Forgetting) for the corresponding vertical position.
Appendix I Extension to More Complex Dataset
Four datasets, including MMLU, MedQA, ARC-C, and CSQA, were combined to construct a more complex dataset, termed MMAC, to evaluate the performance of CNL on more challenging and diverse data distributions. The performance of learning and forgetting during knowledge injection is summarized in Table 9. Results indicate that FT induced 91.7% forgetting during knowledge injection, while CNL achieves zero forgetting.
Table 9: Performance comparison on MMAC. Even on the larger-scale MMAC dataset, FT causes severe forgetting (91.7%), whereas CNL maintains zero forgetting while achieving high learning accuracy.
Method MMAC MMLU MedQA ARC-C CSQA
Learned Forgot Learned Forgot Learned Forgot Learned Forgot Learned Forgot
FT 1878 2949 625 806 665 547 316 819 272 777
(94.9%) (91.7%) (92.2%) (91.9%) (94.5%) (96.3%) (97.5%) (86.6%) (100%) (94.3%)
CNL 1893 0 604 0 700 0 321 0 268 0
(95.7%) (0.0%) (89.1%) (0.0%) (99.4%) (0.0%) (99.1%) (0.0%) (98.5%) (0.0%)

### v2

@@
-Learning the Mechanism of Catastrophic Forgetting: A Perspective from Gradient Similarity
+Collaborative Parameter Learning: Mitigating Forgetting via Parameter-Level Gradient Analysis
 Abstract
-Catastrophic forgetting during knowledge injection severely undermines the continual learning capability of large language models (LLMs).
-While existing methods attempt to mitigate this issue, they often lack a foundational theoretical explanation.
-In this paper, we establish a theoretical framework from the perspective of gradient to explain catastrophic forgetting.
-First, we theoretically prove that strongly negative gradient similarity is the fundamental cause of catastrophic forgetting.
-Subsequently, gradient similarity is employed to identify two types of neurons during catastrophic forgetting: conflicting neurons, which induce forgetting and account for 50%–75% of all neurons, and collaborative neurons, which mitigate forgetting and account for 25%-50%.
-Ultimately, a novel knowledge injection method named Collaborative Neural Learning (CNL) is proposed.
-By freezing conflicting neurons and only updating collaborative neurons, CNL theoretically eliminates catastrophic forgetting under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set.
-Studies on five LLMs, four datasets, and four optimizers prove that CNL achieves zero forgetting under in-set experiments, and reduces the level of catastrophic forgetting by 59.1%–81.7% under out-of-set experiments.
+Catastrophic forgetting during knowledge injection impairs the ability of large language models to acquire new knowledge without overwriting previously mastered knowledge.
+Recent studies analyze forgetting from a gradient-similarity perspective and mitigate forgetting through vector projection.
+However, these methods primarily characterize gradient similarity at the aggregate direction level, leaving the parameter-wise contributions to forgetting underexplored.
+In this paper, we decompose gradient similarity into parameter-wise contributions and identify two types of parameters during forgetting: Conflicting Parameters, whose updates contribute to forgetting and typically account for 50%–75% of parameters, and Collaborative Parameters, whose updates mitigate forgetting and account for 25%–50%.
+Based on this analysis, we propose Collaborative Parameter Learning (CPL), a parameter-wise training rule that freezes Conflicting Parameters and updates only Collaborative Parameters.
+Experiments comparing CPL with seven baseline methods show that CPL learns 20.2%–48.2% more questions with negligible forgetting, while reducing peak VRAM by approximately 3 GB per billion model parameters and computation time by 16.5%.
+Extensive evaluations on parameter consumption, out-of-set generalization, cross-prompt generalization, multimodal tasks, open-ended question answering, and multilingual settings demonstrate that CPL effectively mitigates forgetting across diverse scenarios.
 1 Introduction
-Large language models (LLMs) have shown exceptional knowledge capacity and achieved remarkable potential across a broad range of domains Achiam et al.
-(2023); Touvron et al.
-(2023).
-To continuously expand their knowledge breadth, various knowledge injection methods have been proposed Lewis et al.
-(2020); Wang et al.
-(2024); Ke et al.
-(2023).
-Although these methods enable LLMs to acquire new knowledge, they also cause the forgetting of previously mastered knowledge Zhou et al.
-(2025).
-This phenomenon, known as catastrophic forgetting Korbak et al.
-(2022), severely hinders the continual improvement of LLMs Zhou et al.
-(2024).
-To mitigate catastrophic forgetting, it is crucial to understand its underlying mechanisms.
-Recent studies have attempted to analyze catastrophic forgetting from a “similarity” perspective Bai et al.
-(2024); Wu et al.
-(2024).
-Specifically, they measure the similarity between injected knowledge and previously mastered knowledge using embedding representations Mitchell et al.
-(2022) and knowledge graphs Tan et al.
-(2025); Zhu et al.
-(2025).
-Results show that knowledge highly similar to injected knowledge is more likely to be forgotten Luo et al.
-(2025); Li et al.
-(2024).
-While these studies provide valuable insights into catastrophic forgetting, they fail to rigorously and theoretically reveal its underlying mechanisms.
-Additionally, numerous approaches have been proposed to mitigate catastrophic forgetting, including replay based Huang et al.
-(2024), parameter-efficient fine-tuning based Li et al.
-(2025), and knowledge editing based approaches Ge et al.
-(2024).
-While they alleviate forgetting to some extent, they lack a rigorous theoretical foundation and therefore cannot avoid catastrophic forgetting at theoretically.
-Our work explores the mechanisms of catastrophic forgetting through both theoretical analysis and experimental validation.
-More specifically, we seek to answer the following three key questions regarding catastrophic forgetting:
-$\bullet$ What conditions lead to catastrophic forgetting?
-In Section 3, we demonstrate through rigorous theoretical derivation and extensive empirical results that strongly negativee gradient similarity between injected knowledge and mastered knowledge underlies catastrophic forgetting.
-$\bullet$ Which neurons are responsible for catastrophic forgetting?
-In Section 4, we apply gradient similarity to reveal two distinct types of neurons: conflicting neurons and collaborative neurons.
-Conflicting neurons induce forgetting and comprise 50%–75% of all neurons, whereas collaborative neurons alleviate forgetting and constitute 25%–50%.
-$\bullet$ How to prevent catastrophic forgetting?
-In Section 5, we propose a novel knowledge injection method, CNL, which freezes conflicting neurons and trains collaborative neurons.
-CNL theoretically prevent forgetting under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set.
-Experiments on five models, four datasets, and four optimizers demonstrate CNL effectively mitigates forgetting under both in-set and out-of-set conditions.
-To ensure the reproducibility of this study, all of the codes and datasets are included in Supplementary Material.
-Code and datasets are available at: https://github.com/yangmutian/Collaborative-Neuron-Learning
-Figure 1: Conceptual illustration of our framework. We reveal the mechanism of catastrophic forgetting and propose a novel method to prevent its occurrence from a perspective of gradient similarity.
-2 Definition of Learning and Forgetting from the Perspective of Gradient
+Large language models (LLMs) have shown exceptional knowledge capacity and achieved remarkable potential across a broad range of domains [1, 2].
+To expand their knowledge breadth, various knowledge injection methods have been proposed [3, 4, 5].
+Although these methods enable LLMs to acquire new knowledge, they also cause the forgetting of previously mastered knowledge [6].
+This phenomenon, known as catastrophic forgetting [7], severely undermines the preservation of previously acquired knowledge in LLMs [8].
+Existing literatures on mitigating catastrophic forgetting are typically categorized into four primary paradigms [9].
+Regularization-based methods penalize changes to parameters critical for previously mastered knowledge [10, 11], while parameter-isolation techniques alleviate interference by confining updates to dedicated modules or parameter subsets [12, 13].
+Among these paradigms, gradient-based methods are particularly significant because they formalize forgetting as an optimization-level conflict between injected and mastered knowledge [14, 15].
+By characterizing interference through gradient similarity, these methods mitigate forgetting by projecting the update direction to reduce detrimental interference with mastered knowledge [16, 17, 18].
+However, most existing gradient-based methods treat gradient interference as a single, aggregate vector-level quantity [16, 17, 18], enforcing strict orthogonality between newly injected and previously mastered knowledge to mitigate forgetting.
+As a result, they overlook the parameter-wise contributions to forgetting and impose overly restrictive constraints on knowledge injection.
+Therefore, critical questions arise: do all parameters contribute to forgetting, or only a subset of them?
+How to eliminate the conflicting components between injected and mastered knowledge while preserving the collaborative components, rather than strictly enforcing orthogonality?
+This work attempts to answer these questions as follows:
+$\bullet$ How to characterize catastrophic forgetting from the perspective of gradients?
+In Section 3, we analyze the relationship between gradient similarity and forgetting, revealing that samples with strongly negative gradient similarity are more susceptible to forgetting.
+$\bullet$ Which parameters are responsible for catastrophic forgetting?
+In Section 4, we reveal two distinct types of parameters: Conflicting Parameters and Collaborative Parameters.
+Conflicting Parameters induce forgetting and typically comprise 50%–75% of all parameters, whereas Collaborative Parameters alleviate forgetting and constitute 25%–50%.
+$\bullet$ How to mitigate catastrophic forgetting?
+In Section 5, we propose a novel knowledge injection method, Collaborative Parameter Learning (CPL), which freezes Conflicting Parameters and trains Collaborative Parameters.
+By preserving the collaborative directions between injected and mastered knowledge, rather than enforcing strict orthogonality, CPL learns 20.2%–48.2% more questions with negligible forgetting.
+By replacing high-cost gradient projection with low-cost sign comparison, CPL reduces peak VRAM by 3 GB per billion model parameters and computation time by 16.5%.
+$\bullet$ Is CPL robustness to diverse scenarios?
+In Section 6, we extend CPL to out-of-set generalization, cross-prompt transfer, multimodal tasks, open-ended question answering, and multilingual settings.
+Results indicate that CPL effectively mitigates forgetting in those settings.
+The appendix provides additional experimental details and results.
+Figure 1: Conceptual illustration of our framework. We decompose gradient similarity into parameter-wise contributions, and propose a parameter-wise training rule to mitigate forgetting.
+2 Definition of Learning and Forgetting
 2.1 Learning
-The neural network of LLMs is defined as $f_{\theta}:\mathcal{X}\to\mathcal{Y}$, where the mapping from the input space to the output space is determined by parameters $\theta=(\theta_{1},\dots,\theta_{|\theta|})\in\mathbb{R}^{|\theta|}$.
+The parameters of LLMs are denoted as $\theta=(\theta_{1},\dots,\theta_{|\theta|})\in\mathbb{R}^{|\theta|}$, where $\theta_{j}$ represents the $j$-th parameter.
 Let $\mathcal{I}$ denote a set of knowledge samples to be injected
@@
 Loss on a single injection sample $(x,y)\in\mathcal{I}$ is expressed as $L^{(x,y)}(\theta)=\ell(f_{\theta}(x),y)$, where $\ell(\cdot,\cdot)$ denotes the loss function.
-The total loss on Injection Set $\mathcal{I}$ is expressed as the mean loss over all samples
+The total loss on Injection Set $\mathcal{I}$ is expressed as
 $\mathcal{L}^{\mathcal{I}}(\theta)$ $=\frac{1}{|\mathcal{I}|}\sum_{(x,y)\in\mathcal{I}}L^{(x,y)}(\theta)$
@@
 $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (3)
-Assuming an infinitesimal learning rate $\eta$, the loss after parameter update can be approximated using a first-order Taylor expansion
+The loss after parameter update is approximated using a first-order Taylor expansion
 $\mathcal{L}^{\mathcal{I}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
@@
 $S(\mathcal{I},\mathcal{I})=\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (6)
-Due to the non-negativity of the inner product, we have $\Delta\mathcal{L}^{\mathcal{I}}=-\eta S(\mathcal{I},\mathcal{I})\leq 0$. In practice, parameter updates are rarely zero strictly, leading to
-$\Delta\mathcal{L}^{\mathcal{I}}=-\eta S(\mathcal{I},\mathcal{I})<0$ (7)
-This derivation defines the process of learning: under a sufficiently small learning rate $\eta$, knowledge injection monotonically decreases the loss on Injection Set $\mathcal{I}$.
+Due to the non-negativity of the inner product, we have $\Delta\mathcal{L}^{\mathcal{I}}=-\eta S(\mathcal{I},\mathcal{I})\leq 0$.
+This derivation characterizes knowledge injection: under a sufficiently small learning rate, updating along the negative gradient of $\mathcal{I}$ decreases its loss in the first-order approximation.
 2.2 Forgetting
@@
 $\mathcal{M}=\left\{(x_{1}^{\mathcal{M}},y_{1}^{\mathcal{M}}),\dots,(x_{|\mathcal{M}|}^{\mathcal{M}},y_{|\mathcal{M}|}^{\mathcal{M}})\right\}$ (7)
-The loss on Mastered Set is defined as $\mathcal{L}^{\mathcal{M}}(\theta)$.
-After parameter update according to Eq.
-(3), the loss becomes $\mathcal{L}^{\mathcal{M}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$.
-Assuming an infinitesimal learning rate $\eta$, the expression is expanded through first-order Taylor approximation:
+The loss on $\mathcal{M}$ is defined as $\mathcal{L}^{\mathcal{M}}(\theta)$.
+After parameter update , the loss becomes $\mathcal{L}^{\mathcal{M}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$.
+The expression is expanded through first-order Taylor approximation:
 $\mathcal{L}^{\mathcal{M}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
@@
 $=-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (9)
-We define the Gradient Similarity between Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ as
+We define the gradient similarity between $\mathcal{M}$ and $\mathcal{I}$ as
 $S(\mathcal{M},\mathcal{I})=\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (10)
-Therefore, during knowledge injection, the loss change on Mastered Set is given by
+Therefore, during knowledge injection, the loss change on $\mathcal{M}$ is given by
 $\Delta\mathcal{L}^{\mathcal{M}}=-\eta S(\mathcal{M},\mathcal{I})$ (11)
 Eq.
-(12) reveals a critical insight: the loss on the Mastered Set increases when the gradient similarity is negative.
+(11) reveals a critical insight: the loss on $\mathcal{M}$ increases when the gradient similarity is negative.
 Therefore, mastered samples with more negative gradient similarity (i.e., stronger alignment in the opposite direction) are more susceptible to catastrophic forgetting.
 3 Gradient Similarity Measures Catastrophic Forgetting
-To validate the conclusion of Eq. (12), experiments on five models and four datasets were conducted.
-3.1 Experiment Procedure
-Datasets: We employed standard benchmarks including Massive Multitask Language Understanding (MMLU) Hendrycks et al.
-(2020), Medical Question Answering (MedQA) Jin et al.
-(2021), AI2 Reasoning Challenge (ARC-C) Clark et al.
-(2018), and Commonsense Question Answering (CSQA) Talmor et al.
-(2019).
-The detailed information of datasets is shown in Appendix A.
-Models: The experiments were conducted on Qwen Yang et al. (2024) and LLaMA Grattafiori et al. (2024) families, including Qwen2.5 1.5B, Qwen2.5 3B, Qwen2.5 7B, LLaMA3.2 1B, and LLaMA3.2 3B.
-3.2 Catastrophic Forgetting Happens during Knowledge Injection
-To distinguish Injection Set $\mathcal{I}$ and Mastered Set $\mathcal{M}$, LLMs were asked to answer the questions in datasets (inference details in Appendix B).
-Correctly answered questions were categorized into $\mathcal{M}$, while incorrect ones formed $\mathcal{I}$.
-Appendix C exhibits the distribution of Injection Set $\mathcal{I}$ and Mastered Set $\mathcal{M}$.
-Subsequently, LLMs were trained on Injection Set $\mathcal{I}$ for knowledge injection (training details in Appendix  D).
-The correctly answered samples in Injection Set $\mathcal{I}$ were used to evaluate learning, while the incorrectly answered samples in Mastered Set $\mathcal{M}$ were used to measure forgetting.
-Results in Appendix E reveal that while LLMs successfully learn new knowledge from Injection Set $\mathcal{I}$, they simultaneously suffer forgetting on Mastered Set $\mathcal{M}$.
-Notably, we extended our evaluation to parameter-efficient fine-tuning (PEFT) methods and observed that techniques like Low-Rank Adaptation (LoRA) Hu et al.
-(2022) could not mitigate this catastrophic forgetting.
-Detailed experimental settings and results are provided in Appendix F.
-3.3 Strongly Negative Gradient Similarity Leads to Forgetting
-Ultimately, gradient similarities between samples in Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ were computed.
-In light of Eq.
-(12), samples with negative gradient similarity were ranked by the magnitude of their similarity.
-The top $1/3$ with the largest magnitudes were assigned to the “Sim” group, representing strongly negative similarity, while the bottom $1/3$ were assigned to the “Dissim” group.
-Results in Table 1 indicate that catastrophic forgetting is predominantly concentrated in the “Sim” group.
-This empirical evidence provides strong support for Eq.
-(12), confirming that strongly negative gradient similarity drives catastrophic forgetting.
-Table 1: Relationship between gradient similarity and catastrophic forgetting.
+Table 1: Relationship between gradient similarity and forgetting.
 Results are reported in the format of “Number of forgotten questions (Percentage)”.
-Samples with negative gradient similarity were ranked by magnitude: the top $1/3$ with the largest magnitudes were labeled “Sim” (High Negative Similarity), while the bottom $1/3$ were labeled “Dissim”.
-Results indicate that catastrophic forgetting occurs primarily in the “Sim” group.
+Samples with negative gradient similarity are ranked by magnitude: the top $1/3$ with the largest magnitudes are labeled “Sim”, while the bottom $1/3$ are labeled “Dissim”.
+Results indicate that forgetting occurs primarily in the “Sim” group.
 Model MMLU MedQA ARC-C CSQA
@@
 (0.0%) (33.2%) (11.1%) (12.6%) (0.4%) (25.9%) (0.0%) (13.6%)
-4 Conflicting Neurons Cause Catastrophic Forgetting
-While LLMs exhibit overall forgetting on Mastered Set, it remains unclear whether forgetting is driven by all neurons or only a subset.
-Herein, gradient similarity is employed to identify the neurons contributing to forgetting.
-It should be clarified that in this paper, “neurons” and “parameters” carry essentially the same meaning.
-We use parameters when emphasizing their vector form for computation, and neurons when emphasizing their count.
-The gradient similarity of the $j$-th neuron $\theta_{j}$ between Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ is defined as
-$s_{\theta_{j}}(\mathcal{M},\mathcal{I})=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)$ (13)
-$\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)$ and $\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)$ represent the components of gradients on $\theta_{j}$ for Mastered and Injection Set, respectively.
-The global gradient similarity of LLMs is equivalent to the sum of gradient similarities of individual neurons. Therefore, Eq. (11) is expressed as
-$S(\mathcal{M},\mathcal{I})$ $=\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$
-$=\sum_{j=1}^{|\theta|}\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
-$=\sum_{j=1}^{|\theta|}s_{\theta_{j}}(\mathcal{M},\mathcal{I})$ (14)
-According to Eq.
-(12), if the gradient similarity of neuron $\theta_{j}$ with respect to Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ is negative, this neuron contributes to an increase in loss and induce forgetting.
-We refer to such neurons as Conflicting Neurons.
-The set of conflicting neurons is defined as
-$\theta_{CF}:=\left\{\theta_{j}\mid s_{\theta_{j}}(\mathcal{M},\mathcal{I})<0\right\}$ (15)
-Conversely, if the gradient similarity of neuron $\theta_{j}$ is positive, this neuron contributes to a decrease in loss and mitigate forgetting.
-Such neurons are refered as Collaborative Neurons.
-The set of collaborative neurons is defined as
-$\theta_{CB}:=\left\{\theta_{j}\mid s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0\right\}$ (16)
-Table 2 presents that nearly 25%-50% of neurons contribute to positive gradient similarity and are termed collaborative neurons, while 50%-75% of neurons contribute to negative gradient similarity and are termed conflicting neurons.
-These collaborative neurons are insufficient to counteract the effect of conflicting neurons, thereby leading to the phenomenon of catastrophic forgetting in LLMs.
-This result reveals the coexistence of two types of neurons during catastrophic forgetting, which provide theoretical foundation for the following knowledge injection method.
-Table 2: Distribution of collaborative neurons (COLL) and conflicting neurons (CONF).
-"PROP" represents the proportion of a specific neuron type relative to the total number of neurons.
-“GRAD” represents the gradient similarity sum for that neuron type.
-“TOTAL” represents the gradient similarity sum for all neurons.
-Typically, collaborative neurons account for 25%–50% of all neurons, while conflicting neurons account for 50%–75%.
+3.1 Experiment Procedure
+Datasets: We employ standard benchmarks including Massive Multitask Language Understanding (MMLU) [19], Medical Question Answering (MedQA) [20], AI2 Reasoning Challenge (ARC-C) [21], Commonsense Question Answering (CSQA) [22].
+Additionally, Yahoo Answers [23], AG News [23], and DBPedia [23], NumGLUE-cm [24], NumGLUE-ds [24], C-STANCE [25], and MMBench [26] are used for extensive study.
+Models: The experiments are conducted on Qwen [27] and LLaMA [28] families, including Qwen2.5 1.5B, Qwen2.5 3B, Qwen2.5 7B, LLaMA3.2 1B, and LLaMA3.2 3B.
+The multimodal evaluation is conducted on Qwen2.5vl 3B.
+3.2 Catastrophic Forgetting Happens during Knowledge Injection
+To distinguish Injection Set $\mathcal{I}$ and Mastered Set $\mathcal{M}$, LLMs are asked to answer the questions in datasets (inference details in Appendix B).
+Correctly answered questions are grouped into $\mathcal{M}$, while incorrect ones form $\mathcal{I}$.
+Appendix C exhibits the distribution of $\mathcal{I}$ and $\mathcal{M}$.
+LLMs are trained on $\mathcal{I}$ as detailed in Appendix  D.
+Samples in $\mathcal{I}$ are used to evaluate learning, while samples in $\mathcal{M}$ are used to measure forgetting.
+Table 8 reveals that while LLMs successfully learn new knowledge from $\mathcal{I}$, they simultaneously suffer forgetting on $\mathcal{M}$.
+Similar forgetting is also observed in Low-Rank Adaptation (LoRA) setting [13] as shown in Appendix F.
+3.3 Strongly Negative Gradient Similarity Leads to Forgetting
+Gradient similarity between samples in $\mathcal{M}$ and $\mathcal{I}$ is computed.
+In light of Eq.
+(11), samples with negative gradient similarity are ranked by the magnitude of their similarity.
+The top $1/3$ with the largest magnitudes are assigned to the “Sim” group, representing strongly negative similarity, while the bottom $1/3$ are assigned to the “Dissim” group.
+Table 1 indicates that forgetting is predominantly concentrated in the “Sim” group.
+This evidence provides support for Eq.
+(11), confirming that strongly negative gradient similarity drives forgetting.
+4 Conflicting Parameters Cause Catastrophic Forgetting
+Table 2: Distribution of Collaborative Parameters (COLL) and Conflicting Parameters (CONF).
+“PROP” represents the proportion of a specific parameter.
+“GRAD” represents the gradient similarity sum.
+“TOTAL” represents the gradient similarity sum for all parameters.
+Collaborative Parameters typically account for 25%–50%, while Conflicting Parameters account for 50%–75%.
 Model Metric MMLU MedQA ARC-C CSQA
@@
 Total -2246 -2668 -1427 -1774
-5 Preventing Catastrophic Forgetting by Freezing Conflicting Neurons
-Since conflicting neurons are responsible for catastrophic forgetting, a natural question arises: can catastrophic forgetting be avoided by freezing conflicting neurons and training only collaborative neurons?
-We refer to this learning paradigm as Collaborative Neuron Learning (CNL).
-5.1 Establishment of CNL
-Table 3: Effectiveness of knowledge injection using FT and CNL.
+While LLMs exhibit overall forgetting on $\mathcal{M}$, it remains unclear whether forgetting is driven by all parameters or only a subset.
+To answer this question, gradient similarity of the $j$-th parameter between $\mathcal{M}$ and $\mathcal{I}$ is defined as
+$s_{\theta_{j}}(\mathcal{M},\mathcal{I})=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)$ (12)
+$\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)$ and $\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)$ represent the components of gradients on $\theta_{j}$ for $\mathcal{M}$ and $\mathcal{I}$, respectively.
+The global gradient similarity of LLMs is equivalent to the sum of gradient similarities of individual parameters.
+Therefore, Eq.
+(10) is expressed as
+$S(\mathcal{M},\mathcal{I})$ $=\sum_{j=1}^{|\theta|}\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
+$=\sum_{j=1}^{|\theta|}s_{\theta_{j}}(\mathcal{M},\mathcal{I})$ (13)
+If the gradient similarity of $\theta_{j}$ with respect to $\mathcal{M}$ and $\mathcal{I}$ is negative, this parameter contributes to an increase in loss and induces forgetting.
+We refer to such parameters as Conflicting Parameters.
+Conversely, if the gradient similarity of $\theta_{j}$ is positive, this parameter contributes to a decrease in loss and mitigates forgetting.
+Such parameters are referred to as Collaborative Parameters.
+Table 2 describes the distribution of parameters.
+Approximately 25%–50% of parameters have positive gradient similarity and are termed Collaborative Parameters, while 50%–75% have negative contributions and are termed Conflicting Parameters.
+Moreover, the positive contributions from Collaborative Parameters are generally insufficient to offset the stronger negative contributions from Conflicting Parameters, causing the overall gradient similarity to remain negative.
+5 Mitigating Catastrophic Forgetting by Freezing Conflicting Parameters
+5.1 Establishment of CPL
+Table 3: Performance of knowledge injection using FT and CPL.
 Results are reported in the format of “Number of questions (Percentage)”.
-FT leads to the forgetting of previously mastered knowledge comparable in magnitude to the newly injected knowledge, whereas CNL induces negligible catastrophic forgetting.
+FT leads to significant forgetting of previously mastered knowledge, whereas CPL induces negligible catastrophic forgetting when $\mathcal{M}$ is fully available.
 Model Method MMLU MedQA ARC-C CSQA
@@
 (24.8%) (21.4%) (27.1%) (46.5%) (32.1%) (9.3%) (31.6%) (19.8%)
-CNL 168 0 193 0 163 0 147 0
+CPL 168 0 193 0 163 0 147 0
 (24.8%) (0.0%) (27.4%) (0.0%) (50.3%) (0.0%) (54.0%) (0.0%)
@@
 (19.5%) (13.0%) (18.6%) (25.3%) (33.5%) (27.0%) (31.7%) (12.3%)
-CNL 116 0 119 0 76 0 81 0
+CPL 116 0 119 0 76 0 81 0
 (19.8%) (0.0%) (18.6%) (0.0%) (35.8%) (0.0%) (33.3%) (0.0%)
@@
 (28.0%) (17.5%) (21.9%) (21.9%) (46.3%) (1.7%) (34.5%) (5.3%)
-CNL 189 0 249 0 90 0 144 0
+CPL 189 0 249 0 90 0 144 0
 (41.6%) (0.0%) (49.6%) (0.0%) (67.2%) (0.0%) (73.1%) (0.0%)
@@
 (14.0%) (17.0%) (11.0%) (13.4%) (34.4%) (48.7%) (47.8%) (44.1%)
-CNL 201 0 112 0 335 0 338 0
+CPL 201 0 112 0 335 0 338 0
 (17.9%) (0.0%) (11.8%) (0.0%) (35.1%) (0.0%) (48.5%) (0.0%)
@@
 (18.3%) (11.6%) (22.2%) (13.2%) (24.5%) (8.2%) (19.5%) (4.4%)
-CNL 133 0 125 0 86 0 63 0
+CPL 133 0 125 0 86 0 63 0
 (18.3%) (0.0%) (24.6%) (0.0%) (24.8%) (0.0%) (20.5%) (0.0%)
-To indicate the sign of gradient similarity for the $j$-th neuron, we define an indicator function $I(\cdot)$ as
+Since Conflicting Parameters are responsible for forgetting, a natural question arises: can forgetting be mitigated by freezing Conflicting Parameters and training only Collaborative Parameters?
+We refer to this learning paradigm as Collaborative Parameter Learning (CPL).
+To indicate the sign of gradient similarity for the $j$-th parameter, we define indicator function $I(\cdot)$ as
 $I\!\left(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0\right)=\begin{cases}1,&\text{if }s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0\\ 0,&\text{if }s_{\theta_{j}}(\mathcal{M},\mathcal{I})<0\end{cases}$ (14)
-Applying the indicator function to all neurons yields
-$\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]$ (18)
-$=\big(I(s_{\theta_{1}}(\mathcal{M},\mathcal{I})\geq 0),\ldots,I(s_{\theta_{|\theta|}}(\mathcal{M},\mathcal{I})\geq 0)\big)^{\top}$
-CNL freezes neurons with negative gradient similarity, leading to a parameter update rule defined as
+Applying the indicator function to all parameter yields
+$\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]=\big(I(s_{\theta_{1}}(\mathcal{M},\mathcal{I})\geq 0),\ldots,I(s_{\theta_{|\theta|}}(\mathcal{M},\mathcal{I})\geq 0)\big)^{\top}$ (15)
+CPL freezes parameter with negative gradient similarity, leading to a parameter update rule defined as
 $\theta\leftarrow\theta-\eta\,\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]\odot\nabla_{\theta}L^{\mathcal{I}}(\theta)$ (16)
-where $\odot$ denotes the Hadamard (element-wise) product. According to Eq. (2.2), the change in loss on Mastered Set $\mathcal{M}$ after parameter update is given by
-$\Delta L^{\mathcal{M}}(\theta)$
-$=-\eta\,\nabla_{\theta}L^{\mathcal{M}}(\theta)^{\top}\big(\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]\odot\nabla_{\theta}L^{\mathcal{I}}(\theta)\big)$
+where $\odot$ denotes the Hadamard product. The change in loss on $\mathcal{M}$ after parameter update is given by
+$\Delta L^{\mathcal{M}}(\theta)$ $=-\eta\,\nabla_{\theta}L^{\mathcal{M}}(\theta)^{\top}\big(\mathbb{I}[S(\mathcal{M},\mathcal{I})\geq 0]\odot\nabla_{\theta}L^{\mathcal{I}}(\theta)\big)$
 $=-\eta\sum_{j=1}^{|\theta|}\Big[\nabla_{\theta_{j}}L^{\mathcal{M}}(\theta)\,I(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0)\,\nabla_{\theta_{j}}L^{\mathcal{I}}(\theta)\Big]$
 $=-\eta\sum_{j=1}^{|\theta|}\Big[I(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0)\,s_{\theta_{j}}(\mathcal{M},\mathcal{I})\Big]$ (17)
-In practical settings, it is almost impossible for all parameter gradients to be exactly zero. Therefore, we have
-$\Delta L^{\mathcal{M}}(\theta)$ (21)
-$=-\eta\sum_{j=1}^{|\theta|}\Big[I(s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0)\,s_{\theta_{j}}(\mathcal{M},\mathcal{I})\Big]\;<0$
-This result indicates that, under the assumption of an infinitesimally small learning rate $\eta$, CNL guarantees a monotonic decrease of the loss on Mastered Set $\mathcal{M}$.
-In other words, CNL theoretically prevents catastrophic forgetting.
-To experimentally evaluate the effectiveness of CNL, we performed knowledge injection using both CNL and conventional fine tuning (FT).
-Their forgetting was compared under the constraint that CNL learned at least as many new samples as FT.
-As shown in Table 3, FT forgets a number of mastered samples comparable to the amount of newly injected knowledge, whereas CNL achieve zero forgetting.
-Although CNL exhibits excellent performance in preventing forgetting, we have to point out that CNL achieves zero forgetting only under two restrictive conditions: (i) the learning rate $\eta$ is sufficiently small, and (ii) Mastered Set $\mathcal{M}$ is explicitly known.
-The case where Mastered Set $\mathcal{M}$ is not explicitly available will be discussed in Section 5.4.
-5.2 Extending CNL to More Complex Optimizers
-Table 4: Comparison of knowledge injection performance across different optimizers.
-Qwen2.5 1.5B is selected for experiment.
-Whether using SGD, Momentum, Adam, or AdamW, CNL consistently achieves zero catastrophic forgetting while maintaining effective knowledge injection.
-Optimizer Method MMLU MedQA ARC-C CSQA
+This indicates that, given full access to the samples in $\mathcal{M}$, under a first-order approximation and with an infinitesimally small learning rate, CPL guarantees that the loss on $\mathcal{M}$ does not increase during knowledge injection.
+To experimentally evaluate the effectiveness of CPL, we performed knowledge injection using both CPL and conventional fine tuning (FT) under an ideal situation where $\mathcal{M}$ can be fully accessed.
+The case where $\mathcal{M}$ is not explicitly available will be discussed in Section 6.3.
+CPL and FT are compared under the premise that CPL learns no fewer questions than FT.
+Table 3 shows that FT forgets a number of mastered samples comparable to the amount of newly injected knowledge, whereas CPL forgets zero samples when $\mathcal{M}$ is fully available.
+This demonstrates that CPL effectively mitigates forgetting while maintaining strong capability for knowledge injection.
+5.2 Comparison with Baseline Methods
+Table 4: Comparison with baseline methods on Qwen2.5 1.5B.
+Results are reported in the format of “Number of questions (Percentage)”.
+CPL maintains stronger capability both for knowledge injection and forgetting mitigating.
+MMLU MedQA ARC-C CSQA
 Learned Forgot Learned Forgot Learned Forgot Learned Forgot
-SGD FT 168 188 191 264 104 88 86 163
-(24.8%) (21.4%) (27.1%) (46.5%) (32.1%) (9.3%) (31.6%) (19.8%)
-CNL 168 0 193 0 163 0 147 0
-(24.8%) (0.0%) (27.1%) (0.0%) (50.3%) (0.0%) (54.0%) (0.0%)
-Momentum FT 341 450 257 398 261 169 205 233
-(50.3%) (51.3%) (36.5%) (70.1%) (80.6%) (17.9%) (75.4%) (28.3%)
-CNL 408 0 461 0 291 0 240 0
-(60.2%) (0.0%) (65.5%) (0.0%) (89.8%) (0.0%) (88.2%) (0.0%)
-Adam FT 647 203 704 181 320 143 266 168
-(95.4%) (23.1%) (100.0%) (31.9%) (98.8%) (15.1%) (97.8%) (20.4%)
-CNL 648 0 704 0 323 0 266 0
-(95.6%) (0.0%) (100.0%) (0.0%) (99.7%) (0.0%) (97.8%) (0.0%)
-AdamW FT 638 195 703 174 320 143 251 294
-(94.1%) (22.2%) (99.9%) (30.6%) (98.8%) (15.1%) (92.3%) (35.7%)
-CNL 642 0 703 0 323 0 266 0
-(94.7%) (0.0%) (99.9%) (0.0%) (99.7%) (0.0%) (97.8%) (0.0%)
-After establishing CNL with Stochastic Gradient Descent (SGD) optimizer Rumelhart et al.
-(1985), we further extend CNL to Momentum Rumelhart et al.
-(1985), Adam Kingma (2014), AdamW Loshchilov and Hutter (2019), and arbitrary optimizer.
-In Appendix G, rigorous proofs are provided that the loss on Mastered Set decreases monotonically, as long as the parameter update of Momentum, Adam, AdamW, arbitrary optimizer follows Eq.
-(27), (38), (43), and (58).
-Table 4 demonstrates the experimental results of CNL with difference optimizers.
-When FT is applied, all four optimizers suffer from severe catastrophic forgetting.
-In contrast, CNL effectively prevents catastrophic forgetting across all four optimizers.
-In particular, when using Adam optimizer, CNL is able to successfully inject nearly all (95.6%-100.0%) required knowledge while exhibiting zero forgetting.
-Additionally, the dynamic characteristic of learning and forgetting is analyzed in Appendix.
-H.
-Although CNL freezes conflicting neurons and trains only collaborative neurons, its learning efficiency remains comparable to that of conventional FT.
-This further demonstrates the engineering potential of our CNL.
-5.3 Extending CNL to More Complex Datasets
-To further valid CNL on more complex datasets, MMLU, MedQA, ARC-C, and CSQA were merged into a unified dataset, referred to as MMAC. The performance of FT and CNL on MMAC is studied in Appendix I.
-Results demonstrate that, while successfully injecting 95% knowledge into LLMs, FT leads to 91.7% forgetting of mastered knowledge, while CNL exhibits zero forgetting.
-These findings further demonstrate that CNL can effectively prevent catastrophic forgetting even under more complex and large-scale data conditions.
-5.4 Extending CNL to Out-of-Set Generalization
-We have verified that CNL achieves zero forgetting under in-set setting.
-However, it is hard to explicitly obtain all Mastered Set and Injection Set in practice.
-Therefore, the out-of-set generalization capability of CNL is analyzed.
-First, Injection Set $\mathcal{I}$ was split into two disjoint subsets in a $1:1$ ratio, one used for training and the other reserved for evaluation.
-The “INJECTION" row of Table 5 demonstrates that CNL still achieves zero forgetting because the loss on Mastered Set can be explicitly calculated.
-Subsequently, the same procedure is applied to the Mastered Set $\mathcal{M}$, as reported in the “MASTERED” row of Table 5.
-Under this setting, the loss on the evaluation Mastered subset cannot be directly computed and is instead estimated using the training Mastered subset.
-Due to this estimation error, CNL cannot guarantee a monotonic decrease in the evaluation loss.
-Nevertheless, while zero forgetting cannot be fully achieved, CNL still substantially alleviates catastrophic forgetting, reducing it by 59.1%–81.7% compared to conventional FT.
-Furthermore, the training Mastered subset was combined with Injection Set during training to mitigate forgetting, a strategy known as replay.
-RP follows the same training setup as FT.
-Results in the “RP" row of Table 5 demonstrate that CNL not only learns more new knowledge but also forgets less mastered knowledge compared to replay.
-These results demonstrate that CNL not only achieves zero forgetting under the ideal in-set condition, but also significantly alleviates catastrophic forgetting under more practical out-of-set evaluation settings.
+EWC [10] 164 178 165 209 87 163 97 70
+(24.2%) (20.3%) (23.4%) (36.8%) (26.9%) (17.2%) (35.7%) (8.5%)
+MAS [11] 167 200 184 259 104 80 86 161
+(24.7%) (22.8%) (26.1%) (45.6%) (32.1%) (8.5%) (31.6%) (19.6%)
+MOFO [29] 167 178 193 268 103 92 87 165
+(24.7%) (20.3%) (27.4%) (47.2%) (31.8%) (9.7%) (32.0%) (20.0%)
+WSC [30] 115 97 113 116 76 44 68 96
+(17.0%) (11.0%) (16.1%) (20.4%) (23.5%) (4.7%) (25.0%) (11.7%)
+GPM [18] 165 153 190 257 102 73 83 120
+(24.4%) (17.4%) (27.0%) (45.2%) (31.5%) (7.7%) (30.5%) (14.6%)
+OGD [16] 128 6 142 1 110 3 84 1
+(18.9%) (0.7%) (20.2%) (0.2%) (34.0%) (0.3%) (30.9%) (0.1%)
+A-GEM [17] 129 4 155 1 110 3 84 1
+(19.1%) (0.5%) (22.0%) (0.2%) (34.0%) (0.3%) (30.9%) (0.1%)
+CPL 168 0 193 0 163 0 147 0
+(24.8%) (0.0%) (27.4%) (0.0%) (50.3%) (0.0%) (54.0%) (0.0%)
+Table 5: Comparison of computational cost on Qwen2.5 1.5B.
+CPL replaces the high-precision vector projection with low-precision element-wise sign comparison, reducing peak GPU VRAM by approximately 4.3 GB and time per epoch by 16.5% compared to A-GEM.
+Peak GPU VRAM (G) Time / Epoch (s)
+EWC 29.5 3867
+MAS 31.9 4836
+MOFO 20.1 2170
+WSC 53.0 1808
+GPM 14.9 7504
+OGD 20.6 3738
+A-GEM 20.6 3508
+CPL 16.3 ($\downarrow$ 4.3 GB) 2928 ($\downarrow$ 16.5%)
+We further compare CPL with seven baseline methods: EWC [10], MAS [11], MOFO [29] and WSC [30], GPM [18], OGD [16], and A-GEM [17].
+Table 4 demonstrates that gradient-based methods (OGD and A-GEM) exhibit significantly stronger forgetting mitigation than the other baseline methods.
+Moreover, they are also more conceptually similar to CPL.
+Therefore, we focus on the comparison between CPL and gradient-based approaches.
+In terms of forgetting mitigation, CPL slightly outperforms gradient-based methods, and significantly surpasses other methods.
+In terms of knowledge acquisition, unlike traditional gradient-based methods, which require the gradients of $\mathcal{M}$ and $\mathcal{I}$ to be completely orthogonal, CPL masks only the conflicting component while preserving collaborative component, resulting 20.2%-48.2% more injected samples than OGD and A-GEM within the specified 25 epochs.
+The computational cost is compared in Table 5.
+Gradient-methods require high-cost vector storage for orthogonalization (FP32, accounts for 4 GB per billion parameters), whereas CPL performs element-wise sign comparison and only requires low-costs vector storage (INT8, accounts for 1 GB per billion parameters).
+Therefore, CPL reduces peak GPU VRAM by approximately 3 GB per billion parameters, which is confirmed by the 4.3 GB GPU VRAM reduction observed in Table 5 for Qwen2.5 1.5B.
+Additionally, the simplification from high-cost vector projection to element-wise sign comparison leads to a 16.5% reduction in computational overhead (from 3508 s per epoch for OGD to 2928 s per epoch for CPL).
+5.3 Collaborative Parameters Consumption Analysis
+Since CPL freezes Conflicting Parameters and only updates Collaborative Parameters, it is important to learn that whether Collaborative Parameters are adequately preserved across different datasets.
+To answer this question, the change in Collaborative Parameters for each dataset is recorded after training on a specific dataset.
+Table 6 shows that Collaborative Parameters are task-dependent.
+After training on a dataset, the Collaborative Parameters for that dataset decrease, but the Collaborative Parameters for the other datasets even increase.
+This indicates that CPL does not simply consume a fixed global pool of parameters; instead, the set of Collaborative Parameters changes with the injected knowledge.
+Moreover, CPL is conducted on larger-scale datasets to investigate whether Collaborative Parameters are sufficient for knowledge injection.
+In Appendix G.1, joint multi-dataset learning is conducted.
+Four datasets are mixed together and trained simultaneously.
+Table 10 shows that FT causes 91.7% forgetting on mixed dataset, while CPL forgets 0.0% questions.
+In Appendix G.2, four datasets are trained sequentially to conduct sequential multi-dataset learning.
+Table 11 shows that CPL nearly do not forget mastered questions across all learning orders, while FT suffers significant forgetting.
+Additionally, the learning dynamics of CPL and FT are analyzed in Appendix H.
+Although CPL freezes Conflicting Parameters and only updates Collaborative Parameters, the speed of knowledge injection for CPL is not noticeably slowed than conventionally FT, demonstrating adequate Collaborative Parameters for training.
+6 Extending CPL to Diverse Scenarios
+6.1 Extending CPL to More Optimizers
+In Appendix I, we extend CPL to a broader family of optimizers, including Momentum, Adam, and AdamW.
+Experimental results in Table 12 prove CPL forgets zero questions across different optimizers under the ideal situation that $\mathcal{M}$ is available.
+Moreover, CPL is also theoretically extended to arbitrary optimizers in Appendix I.4.
+6.2 Extending CPL to Various Tasks
+In Appendix J, CPL is further extended to diverse tasks including: sentiment classification (Yahoo Answers, AG News, and DBPedia) [23], open-ended question answering (NumGLUE-cm and NumGLUE-ds) [24], multilingual settings (C-STANCE) [25], and multimodal tasks (MMBench) [26].
+Table 13 demonstrates that CPL effectively mitigates forgetting across diverse tasks.
+6.3 Extending CPL to Out-of-Set Setting
+Table 6: Change in Collaborative Parameters after training.
+The dataset in row is the training dataset, and the dataset in column is the evaluation dataset.
+After training on a specific dataset, its own Collaborative Parameters decrease, but Collaborative Parameters for the others do not decrease.
+DATASET MMLU MedQA ARC-C CSQA
+MMLU -12.1% $20.8\% +$$14.3\% +$$13.3\%$
+MedQA $8.8\% -$8.2% $1.1\% +11$$1.0\%$
+ARC-C $13.9\% +$$17.6\% -$10.5% $14.1\%$
+CSQA $12.4\% +$$23.7\% +$$10.4\% -$9.6%
+In the previous sections, we conducted experiments under an ideal setting that all samples in $\mathcal{M}$ are accessible.
+Herein, CPL is extended to out-of-set generalization settings where only a subset of $\mathcal{M}$ is accessible in Appendix K.
+As shown in Table 14, when $\mathcal{M}$ is partially accessible, although CPL forgets more question than the ideal setting, it still mitigates forgetting by 59.1%–81.7% than conventional FT.
+The result also indicates that incorporating the available subset using CPL is more effective in forgetting mitigation than using replay.
+6.4 Extending CPL to Cross-Prompt Setting
+To study CPL mereley injects specific prompts or the underlying knowledge, four prompts are designed for each question as shown in Appendix L.
+CPL is trained on two prompts and evaluated on the other two prompts.
+Table 15 shows that CPL significantly mitigates forgetting on unseen prompts.
 7 Related Work
-Table 5: Performance evaluation under out-of-set setting.
-Qwen2.5 1.5B is selected for experiment.
-To assess generalization, both Injection Set and Mastered Set were partitioned into training and evaluation subsets.
-This table reports the learning and forgetting performance on unseen evaluation data, demonstrating the excellent generalization capability of CNL.
-Method MMLU MedQA ARC-C CSQA
-Learned Forgot Learned Forgot Learned Forgot Learned Forgot
-Injection FT 24 48 34 75 29 71 22 114
-(7.3%) (5.5%) (9.7%) (13.2%) (17.9%) (7.5%) (16.2%) (13.8%)
-CNL 31 0 39 0 29 0 23 0
-(9.5%) (0.0%) (11.1%) (0.0%) (17.9%) (0.0%) (16.9%) (0.0%)
-Mastered FT 178 110 208 143 104 49 86 94
-(26.3%) (26.0%) (29.5%) (50.4%) (32.1%) (10.4%) (31.6%) (22.8%)
-RP 157 61 124 58 86 14 70 28
-(23.2%) (14.4%) (17.6%) (20.4%) (26.5%) (3.0%) (25.7%) (6.8%)
-CNL 179 46 243 33 116 9 147 18
-(26.4%) (10.9%) (34.5%) (11.6%) (35.8%) (1.9%) (54.0%) (4.4%)
 7.1 Knowledge Injection
-Existing knowledge injection methods are broadly categorized into two types: parametric methods and non-parametric methods Doan et al. (2021); Kotha et al. (2023).
-Parametric methods directly modify model parameters to encode target knowledge via instruction tuning or continued pre-training, but often suffer from catastrophic forgetting of previously acquired knowledge Singhal et al.
-(2023); Han et al.
-(2023); Qiu et al.
-(2024).
-Non-parametric methods, such as retrieval-augmented generation, augment model inference with externally retrieved information without directly modifying model parameters Lewis et al.
-(2020); Ram et al.
-(2023); Xiong et al.
-(2024).
-As a result, they mitigate parameter-level interference, but their effectiveness is constrained by retrieval quality and limited access to implicit knowledge required for complex reasoning Salemi and Zamani (2024); Zhu et al.
-(2024); Kim et al.
-(2025) .
+Knowledge injection methods are commonly divided into parametric and non-parametric approaches [31, 32].
+Parametric methods, such as instruction tuning, update model weights, but can overwrite previously acquired knowledge [33, 34, 35].
+Non-parametric methods, such as retrieval-augmented generation, avoid parameter updates by using external evidence at inference time [3, 36, 37], but remain limited by retrieval quality [38, 39, 40].
 7.2 Mitigating Catastrophic Forgetting
-Approaches to mitigate catastrophic forgetting have been extensively studied De Lange et al. (2021). Existing strategies generally operate along three dimensions.
-Data-based approaches, such as experience replay, mix mastered knowledge with injected knowledge to preserve historical gradient signals, but suffer from storage overhead and privacy concerns.
-Constraint-based approaches, including PEFT, reduce forgetting by restricting the scope of parameter updates or penalizing large deviations from pretrained weights Wang et al.
-(2021); Hu et al.
-(2022).
-Localization-based approaches, such as knowledge editing methods, attempt to identify and modify task-relevant representation regions to achieve targeted knowledge updates with minimal side effects Meng et al.
-(2022, 20222).
-Although effective in practice, most existing methods rely on heuristic designs and lack an explicit characterization of when and why forgetting occurs.
-In contrast, this work seeks to establish a principled connection between forgetting behavior and optimization dynamics.
-6.3 Gradient Analysis in Continual Learning
-From an optimization perspective, the gradient inner product (or gradient similarity) serves as a fundamental indicator of interference between learning objectives Yu et al.
-(2020); Liu et al.
-(2021).
-When gradients induced by new knowledge conflict with those associated with previously learned knowledge, optimization inevitably degrades prior performance, manifesting as catastrophic forgetting Farajtabar et al.
-(2020).
-Previous studies in continual and lifelong learning have leveraged gradient projection or alignment techniques to mitigate such conflicts at a global level Wang et al.
-(2025).
-This work advances this perspective by decomposing gradient inner products at the neuron level, enabling identification of conflicting and collaborative components within the model.
-By selectively suppressing neuron-level conflicting gradients while preserving collaborative ones, forgetting mitigation is transformed from heuristic data balancing into a theoretically grounded training rule that is compatible with standard optimizers.
+For model adaptation, forgetting is typically mitigated through replay, constraints, localization, or update selection [9].
+Replay mixes old and new data but introduces storage and privacy costs [41]; constraint methods such as PEFT, EWC, and MAS restrict updates or protect important weights [12, 13, 10, 11]; localization and editing methods target task-relevant regions [42, 43]; and recent methods such as MOFO and WSC select or consolidate updates during adaptation [29, 30].
+Gradient similarity provides a direct signal of interference between learning objectives [14, 15]: conflicting gradients can degrade prior knowledge and induce forgetting [16].
+Prior projection or alignment methods, including OGD, A-GEM, and GPM, mitigate such conflicts at a global level [16, 17, 18, 44].
+Limitations
+First, our analysis relies on a first-order approximation.
+However, LLMs are generally non-convex.
+Future work could explore non-linear dynamics of forgetting.
+Second, CPL assumes access to $\mathcal{M}$.
+Although we show that CPL can still mitigate forgetting when only a subset of $\mathcal{M}$ is accessible, it remains an open question how to best estimate gradient similarity without direct access to $\mathcal{M}$.
+Future work could explore proxy methods for gradient estimation.
+Third, the gradient computation of $\mathcal{M}$ requires additional computational cost.
+Although CPL reduces computational cost compared with other existing gradient-based methods, it still incurs more cost than conventional FT.
+Future work could explore more efficient approximations of gradient similarity.
 Conclusion
-Catastrophic forgetting during knowledge injection poses a fundamental challenge to the continual learning capability of LLMs.
-In this work, we establish a gradient-based theoretical framework to explain the underlying mechanism of catastrophic forgetting.
-We show that strongly negative gradient similarity is the fundamental cause of forgetting and further leverage this insight to categorize neurons into two types: conflicting neurons, which induce forgetting and account for 50%–75% of all neurons, and collaborative neurons, which mitigate forgetting and account for the remaining 25%–50%.
-Based on this analysis, we propose a novel knowledge injection method, CNL, which freezes conflicting neurons and updates only collaborative neurons.
-We theoretically show that, under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set, CNL prevents catastrophic forgetting.
-Extensive experiments on five LLMs, four datasets, and four optimizers further validate our analysis.
-CNL achieves zero forgetting under in-set evaluation and reduces catastrophic forgetting by 59.1%–81.7% under more practical out-of-set settings.
-Overall, this work provides both theoretical insights and practical solutions for mitigating catastrophic forgetting in knowledge injection, offering a principled foundation for continual knowledge acquisition in LLMs.
-Acknowledgments
-This work was supported by the Noncommunicable Chronic Diseases-National Science and Technology Major Project under Grants 2024ZD0522702 and 2024ZD0528306, and the National Natural Science Foundation of China under Grant 62433007.
+Catastrophic forgetting during knowledge injection prevents LLMs from retaining previously learned knowledge.
+We show that strongly negative gradient similarity drives forgetting, and categorize parameters into Conflicting and Collaborative Parameters.
+CPL freezes Conflicting Parameters and updates only Collaborative Parameters.
+Experiments show that CPL significantly mitigates forgetting while maintaining knowledge injection capability across diverse scenarios.
+Impact Statement
+This study provides parameter-level understanding for forgetting during knowledge injection and reveals the role of Conflicting and Collaborative parameters.
+Building on this insight, CPL is proposed by freezing Conflicting Parameters and updating only Collaborative Parameters.
+Like other knowledge injection methods, CPL may inject incorrect, biased, or harmful knowledge if the injected knowledge is poorly curated, so practical deployment requires careful data governance.
 Appendix A Dataset Details
-To verify our conclusions, we selected four datasets for experimentation: MMLU, MedQA, ARC-C, and CSQA. The data processing method is described as follows:
-  • MMLU: 1/10 subsampling is performed to result in 1, 555 remaining questions.
-  • MedQA: 1/10 subsampling is performed to result in 1, 273 remaining questions.
-  • ARC-C: 1/2 subsampling is performed to result in 1, 270 remaining questions.
-  • CSQA: 1/10 subsampling is performed to result in 1, 096 remaining questions.
+MMLU, MedQA, ARC-C, and CSQA are employed as foundation datasets, which are 1/10, 1/10, 1/2, and 1/10 randomly subsampled from the original datasets, respectively, resulting in 1,555, 1,273, 1,270, and 1,096 questions, respectively.
+Yahoo Answers, AG News, DBPedia, NumGLUE-cm, NumGLUE-ds, C-STANCE, and MMBench are used for extensive study.
+All datasets are randomly subsampled to 500 questions each, except for MMBench, which contains 500 Chinese and 500 English questions.
 Appendix B Model Inference Details
-Model inference was conducted in a Python (3.12.2) environment, with the main packages including transformers (4.48.0) and torch (2.2.0).
-To ensure experimental reproducibility, the temperature was set to 0.
-Model inference precision was set to FP32, and experiments were conducted on a single A800 GPU.
+Model inference is conducted in a Python (3.12.2) environment, with the main packages including transformers (4.48.0) and torch (2.2.0).
+To ensure experimental reproducibility, the temperature is set to 0.
+Model inference precision is set to FP32, and experiments are conducted on a single A800 GPU.
 Appendix C Model Inference Results
-To identify Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$, LLMs were asked to answer the questions in the datasets.
-Correctly answered questions were assigned to Mastered Set $\mathcal{M}$, while incorrectly answered questions were assigned to Injection Set $\mathcal{I}$.
+To identify Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$, LLMs are asked to answer the questions in the datasets.
+Correctly answered questions are assigned to Mastered Set $\mathcal{M}$, while incorrectly answered questions are assigned to Injection Set $\mathcal{I}$.
 The distribution of different sets are described in Table.
-6.
+7.
 Table 7: Number of questions in Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$.
 “INJECTION” denotes the number of samples in Injection Set (answered incorrectly), while “MASTERED” denotes the number of samples in Mastered Set (answered correctly).
-Model MMLU MedQA ARC-C CSQA
+MMLU MedQA ARC-C CSQA
 Injection Mastered Injection Mastered Injection Mastered Injection Mastered
@@
 Appendix D Model Training Details
-The environment of model training was essentially the same as that for model inference.
-The number of epochs was set to 25.
+The environment of model training is essentially the same as that for model inference.
+The number of epochs is set to 25.
 For the SGD optimizer, the learning rate included three values according to their magnitude of gradient: $2\text{e-}9$, $2\text{e-}7$, and $1\text{e-}7$ (see the code for exact configurations).
-For Momentum optimizer, the learning rate was consistent with SGD, and the momentum coefficient was set to 0.9.
-For Adam optimizer (without regularization), $\beta_{1}$ and $\beta_{2}$ were 0.9 and 0.999 respectively, and $\epsilon$ was $1\text{e-}8$.
-For AdamW optimizer, the regularization parameter (weight decay) was $1\text{e-}1$.
-Model training was performed in FP32 precision on a single NVIDIA A800 GPU.
+For Momentum optimizer, the learning rate is consistent with SGD, and the momentum coefficient is set to 0.9.
+For Adam optimizer (without regularization), $\beta_{1}$ and $\beta_{2}$ are 0.9 and 0.999 respectively, and $\epsilon$ is $1\text{e-}8$.
+For AdamW optimizer, the regularization parameter (weight decay) is $1\text{e-}1$.
+Model training is performed on a single NVIDIA A800 GPU.
 Appendix E Model Training Results
-The samples in Injection Set $\mathcal{I}$ were trained for knowledge injection.
-After training, the number of correctly answered questions in Injection Set $\mathcal{I}$ was employed to evaluate learning, while the number of incorrectly answered questions in Mastered Set $\mathcal{M}$ was employed to evaluate forgetting.
+The samples in Injection Set $\mathcal{I}$ are trained for knowledge injection.
+After training, the number of correctly answered questions in Injection Set $\mathcal{I}$ is employed to evaluate learning, while the number of incorrectly answered questions in Mastered Set $\mathcal{M}$ is employed to evaluate forgetting.
 The results in Table.
@@
 After training, although new knowledge is successfully injected into LLMs, severe catastrophic forgetting is also observed.
-Model MMLU MedQA ARC-C CSQA
+MMLU MedQA ARC-C CSQA
 Learned Forgot Learned Forgot Learned Forgot Learned Forgot
@@
 Appendix F Evaluation on Parameter-Efficient Fine-Tuning
-We evaluate LoRA as a representative parameter-efficient fine-tuning (PEFT) baseline.
-All experiments are conducted on the Qwen2.5-1.5B model.
-For a fair comparison with standard fine-tuning (Appendix D), we adopt the same training protocol, using the SGD optimizer with a learning rate of $\eta=1\times 10^{-5}$.
+We evaluate LoRA as a representative parameter-efficient fine-tuning baseline.
+All experiments are conducted on the Qwen2.5-1.5B.
+For a fair comparison with FT, we adopt the same training protocol with Appendix D, using the SGD optimizer with a learning rate of $\eta=1\times 10^{-5}$.
 The LoRA configuration employs a rank of $r=16$ and a scaling factor of $\alpha=32$, with LoRA modules applied to all linear layers in the attention mechanism.
 The batch size and number of training epochs are identical to those used in Appendix D.
-As shown in Table 8, LoRA achieves learning performance comparable to full fine-tuning, but does not alleviate catastrophic forgetting, exhibiting similarly severe forgetting across all evaluated datasets.
-Table 8: Comparison of FT and LoRA under during knowledge injection. Forgetting cannot be effectively prevented when LoRA is employed.
-Method MMLU MedQA ARC-C CSQA
+As shown in Table 9, when LoRA achieves learning performance comparable to FT, LoRA does not alleviate catastrophic forgetting, exhibiting similarly severe forgetting across all evaluated datasets.
+Table 9: Comparison of FT and LoRA during knowledge injection. Forgetting cannot be effectively mitigated when LoRA is employed.
+MMLU MedQA ARC-C CSQA
 Learned Forgot Learned Forgot Learned Forgot Learned Forgot
@@
 (24.8%) (20.1%) (27.4%) (50.4%) (32.1%) (8.4%) (31.6%) (19.9%)
-Appendix G Theoretical Derivation for More Complex Optimizers
-G.1 Momentum Optimizer
-When performing gradient descent with learning rate $\eta$, the update rule for parameters $\theta$ in the Momentum optimizer is as follows:
-$m$ $\leftarrow\beta m+\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (22)
-$\theta$ $\leftarrow\theta-\eta m$
-where $m$ performs an exponential weighted average of historical gradients for accelerated convergence.
-$\beta\in[0,1)$ is the momentum coefficient, used to control the decay rate of historical momentum information.
-The impact of parameter updates on existing knowledge is analyzed. Similar to Eq. (2.2), we perform a Taylor expansion on loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over Mastered Set
-$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta m)$ (23)
-$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta m)$
-The change in loss on Mastered Set is:
-$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}m}_{S^{mom}(\mathcal{M},\mathcal{I})}$ (24)
-Here, $S_{mom}(\mathcal{M},\mathcal{I})$ is defined as the momentum similarity.
-If $S_{mom}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting.
-Utilizing the additivity of inner product, we decompose the global momentum similarity onto each neuron
-$S^{mom}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{mom}$ (25)
-where $s_{\theta_{j}}^{mom}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot m_{j}$. Based on the sign of $s_{\theta_{j}}^{mom}$, we classify the neurons into two categories:
-$\theta_{CF}$ $=\left\{\theta_{j}\mid s_{\theta_{j}}^{mom}(\mathcal{M},\mathcal{I})<0\right\}$ (Conflicting Neurons) (26)
-$\theta_{CO}$ $=\left\{\theta_{j}\mid s_{\theta_{j}}^{mom}(\mathcal{M},\mathcal{I})\geq 0\right\}$ (Collaborative Neurons)
-Similar to CNL with SGD, the parameter update rule of CNL for momentum is
-$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{mom}\geq 0)\odot m\right]$ (27)
-Under this rule, the loss in Master Set is
-$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{mom}\geq 0)\odot m\right]$ (28)
-$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{mom}\geq 0)\cdot(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot m_{j})\right]$
-$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{mom}\geq 0)\cdot s_{\theta_{j}}^{mom}\right]\leq 0$
-In practical settings, it is nearly impossible for all gradients to be exactly zero. Therefore we get
-$\Delta\mathcal{L}^{\mathcal{M}}<0$ (29)
-G.2 Adam Optimizer (without weight decay)
-The conventional update rule for the Adam optimizer is as follows:
-$m$ $\leftarrow\beta_{1}m+(1-\beta_{1})g_{t}$ (30)
-$v$ $\leftarrow\beta_{2}v+(1-\beta_{2})g_{t}^{2}\quad$
-$\hat{m}$ $=\frac{m}{1-\beta_{1}^{t}},\quad\hat{v}=\frac{v}{1-\beta_{2}^{t}}\quad(\text{bias correction})$
-$\theta$ $\leftarrow\theta-\eta\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}$
-where $g_{t}=\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ represents the gradient vector at step $t$.
-$m$ and $v$ are the exponential moving average estimates of the first moment (mean) and the second raw moment (uncentered variance) of the gradients, respectively.
-$\beta_{1},\beta_{2}\in[0,1)$ are the exponential decay rates for the first and second moment estimates, respectively.
-$\hat{m}$ and $\hat{v}$ are the bias-corrected moment estimates, intended to counteract the bias caused by initializing $m,v$ to zero during the initial training phase.
-$\epsilon$ is a numerical stability constant added to prevent division by zero.
-For notational convenience, we denote the actual update direction calculated by Adam as the vector $U_{adam}$
-$U^{adam}=\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}$ (31)
-Therefore, the conventional update rule for the Adam optimizer is
-$\theta=\theta-\eta U^{adam}.$ (32)
-Next, we analyze the impact of parameter updates on existing knowledge. We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set:
-$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta U^{adam})$ (33)
-$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta U^{adam})$
-The change in loss on the mastered set is:
-$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}U^{adam}}_{S^{adam}(\mathcal{M},\mathcal{I})}$ (34)
-Here, $S_{adam}(\mathcal{M},\mathcal{I})$ is defined as the Adam similarity. If $S_{adam}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting.
-Utilizing the additivity of the inner product, we decompose the global Adam similarity onto each neuron:
-$S^{adam}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{adam}$ (35)
-where the similarity component of the $j$-th neuron is:
-$s_{\theta_{j}}^{adam}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adam}_{j}$ (36)
-We classify the neurons into two categories:
-$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adam}<0\}\quad\text{(Conflicting Neurons)}$ (37)
-$\theta_{CO}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adam}\geq 0\}\quad\text{(collaborative Neurons)}$
-The modified CNL parameter update formula is:
-$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{adam}\geq 0)\odot U^{adam}\right]$ (38)
-Substituting the CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
-$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{adam}\geq 0)\odot U^{adam}\right]$ (39)
-$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{adam}\geq 0)\cdot\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adam}_{j}\right)\right]$
-$=-\eta\sum_{j=1}^{|\theta|}[I(s_{\theta_{j}}^{adam}\geq 0)\cdot s_{\theta_{j}}^{adam}]\leq 0$
-In practical engineering,
-$\Delta\mathcal{L}^{\mathcal{M}}<0$ (40)
-G.3 AdamW Optimizer
-The AdamW optimizer introduces decoupled weight decay on top of Adam, separating the weight decay term from the gradient update. Its update rule is as follows:
-$m$ $\leftarrow\beta_{1}m+(1-\beta_{1})g_{t}$ (41)
-$v$ $\leftarrow\beta_{2}v+(1-\beta_{2})g_{t}^{2}$
-$\hat{m}$ $=\frac{m}{1-\beta_{1}^{t}},\quad\hat{v}=\frac{v}{1-\beta_{2}^{t}}$
-$\theta$ $\leftarrow\theta-\eta\left(\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}+\lambda\theta\right)$
-where the definitions of $m,v,\hat{m},\hat{v},\beta_{1},\beta_{2},\epsilon$ are consistent with the Adam optimizer.
-$\lambda$ represents the Weight Decay Coefficient.
-Note that the update vector $U_{adamw}$ here includes not only the gradient-based adaptive step size but also the regularization term $\lambda\theta$ pointing towards the origin.
-For notational convenience, we denote the actual total update direction calculated by AdamW (including weight decay) as the vector $U^{adamw}$:
-$U^{adamw}=\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}+\lambda\theta$ (42)
-The rule of parameter update becomes
-$\theta=\theta-\eta U^{adamw}.$ (43)
-Next, we analyze the impact of parameter updates on existing knowledge. We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set:
-$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta U^{adamw})$ (44)
-$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta U^{adamw})$
-The change in loss on Mastered Set is:
-$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}U^{adamw}}_{S^{adamw}(\mathcal{M},\mathcal{I})}$ (45)
-Here, $S_{adamw}(\mathcal{M},\mathcal{I})$ is defined as the AdamW similarity. If $S_{adamw}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting.
-Utilizing the additivity of the inner product, we decompose the global AdamW similarity onto each neuron
-$S^{adamw}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{adamw}$ (46)
-where the similarity component of the $j$-th neuron is:
-$s_{\theta_{j}}^{adamw}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adamw}_{j}$ (47)
-We classify the neurons into two categories:
-$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adamw}<0\}\quad\text{(Conflicting Neurons)}$ (48)
-$\theta_{CB}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adamw}\geq 0\}\quad\text{(collaborative Neurons)}$
-The modified CNL parameter update formula is
-$\theta\leftarrow\theta-\eta\left[I(S^{adamw}\geq 0)\odot U^{adamw}\right]$ (49)
-Substituting the CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
-$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\cdot}\left[\mathbb{I}(\mathbf{S}^{adamw}\geq 0)\odot U^{adamw}\right]$ (50)
-$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{adamw}\geq 0)\cdot\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adamw}_{j}\right)\right]$
-$=-\eta\sum_{j=1}^{|\theta|}[I(s_{\theta_{j}}^{adamw}\geq 0)\cdot s_{\theta_{j}}^{adamw}]\leq 0$
-In practical engineering
-$\Delta\mathcal{L}^{\mathcal{M}}<0$ (51)
-G.4 General Optimizers
-To derive a unified theoretical framework, we abstract the specific update rules of different optimizers into a generalized form.
-Regardless of whether the optimizer utilizes momentum, adaptive learning rates, or weight decay, its parameter update can ultimately be expressed as applying a specific update vector $F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ to the current parameters.
-The general update rule is defined as:
-$\theta$ $\leftarrow\theta-\eta F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ (52)
-where $\eta$ is the global learning rate.
-Next, we analyze the impact of this general parameter update on existing knowledge.
-We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set
-$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)))$ (53)
-$\approx\mathcal{L}^{\mathcal{M}}(\theta)-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$
-The change in loss on the mastered set is:
-$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))}_{S^{opt}(\mathcal{M},\mathcal{I})}$ (54)
-Utilizing the additivity of the inner product, we decompose the global similarity onto each neuron $j$:
-$S^{opt}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{opt}$ (55)
-where the similarity component of the $j$-th neuron is determined by the projection of the optimizer’s update scalar:
-$s_{\theta_{j}}^{opt}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot F^{opt}_{j}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ (56)
-Based on the sign of $s_{\theta_{j}}^{opt}$, we universally classify the neurons into two categories for any optimizer:
-$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{opt}<0\}\quad\text{(Conflicting Neurons)}$ (57)
-$\theta_{CB}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{opt}\geq 0\}\quad\text{(Collaborative Neurons)}$
-The generalized CNL parameter update formula is:
-$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{opt}\geq 0)\odot F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$ (58)
-Substituting the generalized CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
-$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{opt}\geq 0)\odot F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$ (59)
-$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{opt}\geq 0)\cdot(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot F^{opt}_{j}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$
-$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{opt}\geq 0)\cdot s_{\theta_{j}}^{opt}\right]\leq 0$
-In practical engineering, we have:
-$\Delta\mathcal{L}^{\mathcal{M}}<0$ (60)
-Appendix H Dynamic Characteristic of CNL
+Appendix G Extending CPL to Larger-Scale Datasets
+G.1 Extending CPL to Joint Multi-Dataset Learning
+Four datasets are combined to construct a more complex dataset, termed MMAC, to evaluate the performance of CPL on larger-scale datasets.
+Table 10 indicates that FT induced 91.7% forgetting during knowledge injection, while CPL forgets zero questions under the ideal situation that Mastered Set $\mathcal{M}$ is fully accessed.
+Table 10: Performance of joint multi-dataset learning. On the larger-scale MMAC dataset, FT causes severe forgetting (91.7%), whereas CPL maintains 0.0% forgetting.
+MMAC MMLU MedQA ARC-C CSQA
+Learned Forgot Learned Forgot Learned Forgot Learned Forgot Learned Forgot
+FT 1878 2949 625 806 665 547 316 819 272 777
+(94.9%) (91.7%) (92.2%) (91.9%) (94.5%) (96.3%) (97.5%) (86.6%) (100%) (94.3%)
+CPL 1893 0 604 0 700 0 321 0 268 0
+(95.7%) (0.0%) (89.1%) (0.0%) (99.4%) (0.0%) (99.1%) (0.0%) (98.5%) (0.0%)
+G.2 Extending CPL to Sequential Multi-Dataset Learning
+The datasets are trained sequentially according to the “Order” column in Table 11. CPL consistently mitigates catastrophic forgetting across different task orders.
+Table 11: Performance of sequential multi-dataset learning.
+AR, CS, $M$, and Me denote ARC-C, CSQA, MMLU, and MedQA, respectively.
+Results indicate that CPL nearly do not forget questions across all orders while maintaining effective learning, under the ideal situtation that Mastered Set $\mathcal{M}$ is fully accessed.
+Order MMLU MedQA ARC-C CSQA
+Learned Forgot Learned Forgot Learned Forgot Learned Forgot
+AR-CS-$M$-Me 151 0 167 0 109 0 90 0
+$M$-Me-AR-CS 141 1 149 0 113 0 94 0
+CS-AR-Me-$M$ 150 0 148 0 106 0 98 0
+Me-AR-CS-$M$ 145 0 152 0 110 0 92 0
+Appendix H Dynamic Characteristic of CPL
 The dynamic characteristics of learning and forgetting during knowledge injection are shown in Fig.
 2.
-Although CNL freezes conflicting neurons, its training efficiency does not significantly lag behind that of conventional FT.
+Although CPL freezes Conflicting Parameters and only updates Collaborative Parameters, the speed of knowledge injection for CPL is not noticeably slowed than FT, This indicates that Collaborative Parameters are sufficient for training, and the knowledge injection capability of CPL is not compromised by freezing Conflicting Parameters.
 MMLU
@@
 The inner labels indicate the metric type (Learning vs Forgetting) for the corresponding vertical position.
-Appendix I Extension to More Complex Dataset
-Four datasets, including MMLU, MedQA, ARC-C, and CSQA, were combined to construct a more complex dataset, termed MMAC, to evaluate the performance of CNL on more challenging and diverse data distributions.
-The performance of learning and forgetting during knowledge injection is summarized in Table 9.
-Results indicate that FT induced 91.7% forgetting during knowledge injection, while CNL achieves zero forgetting.
-Table 9: Performance comparison on MMAC. Even on the larger-scale MMAC dataset, FT causes severe forgetting (91.7%), whereas CNL maintains zero forgetting while achieving high learning accuracy.
-Method MMAC MMLU MedQA ARC-C CSQA
-Learned Forgot Learned Forgot Learned Forgot Learned Forgot Learned Forgot
-FT 1878 2949 625 806 665 547 316 819 272 777
-(94.9%) (91.7%) (92.2%) (91.9%) (94.5%) (96.3%) (97.5%) (86.6%) (100%) (94.3%)
-CNL 1893 0 604 0 700 0 321 0 268 0
-(95.7%) (0.0%) (89.1%) (0.0%) (99.4%) (0.0%) (99.1%) (0.0%) (98.5%) (0.0%)
+Appendix I Extending CPL to More Optimizers
+I.1 Momentum Optimizer
+When performing gradient descent with learning rate $\eta$, the update rule for parameters $\theta$ in the Momentum optimizer is as follows:
+$m$ $\leftarrow\beta m+\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ (18)
+$\theta$ $\leftarrow\theta-\eta m$
+where $m$ performs an exponential weighted average of historical gradients for accelerated convergence. Similar to Eq. (2.2), the first-order change in loss on Mastered Set is
+$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}m}_{S^{mom}(\mathcal{M},\mathcal{I})}.$ (19)
+Here, $S^{mom}(\mathcal{M},\mathcal{I})$ is defined as the momentum similarity. Utilizing the additivity of inner product, we decompose the global momentum similarity onto each neuron:
+$S^{mom}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{mom},\quad s_{\theta_{j}}^{mom}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot m_{j}.$ (20)
+The parameter update rule of CPL for momentum is
+$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{mom}\geq 0)\odot m\right].$ (21)
+Under this rule, the first-order loss change on Mastered Set is
+$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\sum_{j=1}^{|\theta|}I(s_{\theta_{j}}^{mom}\geq 0)\cdot s_{\theta_{j}}^{mom}\leq 0.$ (22)
+I.2 Adam Optimizer (without weight decay)
+The conventional update rule for the Adam optimizer is as follows:
+$m$ $\leftarrow\beta_{1}m+(1-\beta_{1})g_{t}$ (23)
+$v$ $\leftarrow\beta_{2}v+(1-\beta_{2})g_{t}^{2}$
+$\hat{m}$ $=\frac{m}{1-\beta_{1}^{t}},\quad\hat{v}=\frac{v}{1-\beta_{2}^{t}}$
+$\theta$ $\leftarrow\theta-\eta\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}$
+where $g_{t}=\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$. Denote the actual Adam update direction as
+$U^{adam}=\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}.$ (24)
+The first-order change in loss on Mastered Set is
+$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}U^{adam}}_{S^{adam}(\mathcal{M},\mathcal{I})}.$ (25)
+We decompose the global Adam similarity as
+$S^{adam}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{adam},\quad s_{\theta_{j}}^{adam}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adam}_{j}.$ (26)
+The CPL update rule for Adam is
+$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{adam}\geq 0)\odot U^{adam}\right],$ (27)
+which yields
+$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\sum_{j=1}^{|\theta|}I(s_{\theta_{j}}^{adam}\geq 0)\cdot s_{\theta_{j}}^{adam}\leq 0.$ (28)
+I.3 AdamW Optimizer
+The AdamW optimizer introduces decoupled weight decay on top of Adam:
+$m$ $\leftarrow\beta_{1}m+(1-\beta_{1})g_{t}$ (29)
+$v$ $\leftarrow\beta_{2}v+(1-\beta_{2})g_{t}^{2}$
+$\hat{m}$ $=\frac{m}{1-\beta_{1}^{t}},\quad\hat{v}=\frac{v}{1-\beta_{2}^{t}}$
+$\theta$ $\leftarrow\theta-\eta\left(\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}+\lambda\theta\right).$
+For notational convenience, we denote the actual total update direction calculated by AdamW as
+$U^{adamw}=\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}+\lambda\theta.$ (30)
+The first-order loss change is
+$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}U^{adamw}}_{S^{adamw}(\mathcal{M},\mathcal{I})}.$ (31)
+We decompose it as
+$S^{adamw}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{adamw},\quad s_{\theta_{j}}^{adamw}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adamw}_{j}.$ (32)
+The CPL update rule for AdamW is
+$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{adamw}\geq 0)\odot U^{adamw}\right],$ (33)
+and the first-order loss change satisfies
+$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\sum_{j=1}^{|\theta|}I(s_{\theta_{j}}^{adamw}\geq 0)\cdot s_{\theta_{j}}^{adamw}\leq 0.$ (34)
+I.4 Arbitrary Optimizers
+To derive a unified first-order framework, we abstract the specific update rules of different optimizers into a generalized form.
+Regardless of whether the optimizer utilizes momentum, adaptive learning rates, or weight decay, its parameter update can ultimately be expressed as applying a specific update vector $F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ to the current parameters:
+$\theta\leftarrow\theta-\eta F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)).$ (35)
+The first-order loss change on the mastered set is
+$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))}_{S^{opt}(\mathcal{M},\mathcal{I})}.$ (36)
+Utilizing the additivity of the inner product, we decompose the global similarity onto each neuron:
+$S^{opt}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{opt},\quad s_{\theta_{j}}^{opt}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot F^{opt}_{j}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)).$ (37)
+The generalized CPL parameter update formula is
+$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{opt}\geq 0)\odot F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right].$ (38)
+Substituting the generalized CPL update rule back into the first-order approximation gives
+$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\sum_{j=1}^{|\theta|}I(s_{\theta_{j}}^{opt}\geq 0)\cdot s_{\theta_{j}}^{opt}\leq 0.$ (39)
+I.5 Experimental Results
+When applying CPL to Momentum, Adam, and AdamW optimizers, we observe consistent performance in mitigating catastrophic forgetting across all four datasets.
+As shown in Table 12, regardless of the optimizer used, CPL achieves zero forgetting while maintaining effective knowledge injection, demonstrating its general applicability across different optimization.
+Table 12: Comparison of knowledge injection performance across different optimizers.
+Qwen2.5 1.5B is selected for experiment.
+Whether using SGD, Momentum, Adam, or AdamW, CPL consistently forgets zero question when Mastered Set $\mathcal{M}$ is fully accessed.
+Optimizer Method MMLU MedQA ARC-C CSQA
+Learned Forgot Learned Forgot Learned Forgot Learned Forgot
+SGD FT 168 188 191 264 104 88 86 163
+(24.8%) (21.4%) (27.1%) (46.5%) (32.1%) (9.3%) (31.6%) (19.8%)
+CPL 168 0 193 0 163 0 147 0
+(24.8%) (0.0%) (27.1%) (0.0%) (50.3%) (0.0%) (54.0%) (0.0%)
+Momentum FT 341 450 257 398 261 169 205 233
+(50.3%) (51.3%) (36.5%) (70.1%) (80.6%) (17.9%) (75.4%) (28.3%)
+CPL 408 0 461 0 291 0 240 0
+(60.2%) (0.0%) (65.5%) (0.0%) (89.8%) (0.0%) (88.2%) (0.0%)
+Adam FT 647 203 704 181 320 143 266 168
+(95.4%) (23.1%) (100.0%) (31.9%) (98.8%) (15.1%) (97.8%) (20.4%)
+CPL 648 0 704 0 323 0 266 0
+(95.6%) (0.0%) (100.0%) (0.0%) (99.7%) (0.0%) (97.8%) (0.0%)
+AdamW FT 638 195 703 174 320 143 251 294
+(94.1%) (22.2%) (99.9%) (30.6%) (98.8%) (15.1%) (92.3%) (35.7%)
+CPL 642 0 703 0 323 0 266 0
+(94.7%) (0.0%) (99.9%) (0.0%) (99.7%) (0.0%) (97.8%) (0.0%)
+Appendix J Extending CPL to Various Tasks
+CPL is extended to various tasks including: sentiment classification (Yahoo Answers, AG News, and DBPedia) [23], open-ended question answering (NumGLUE-cm and NumGLUE-ds) [24], multilingual settings (C-STANCE) [25], and multimodal tasks (MMBench) [26].
+In Table 13, each cell is reported as “Learned/Forgotten”.
+CPL keeps the forgotten count much smaller across both tasks compared with FT, demonstrating its effectiveness across a wide range of tasks.
+Table 13: Learning results on 7 tasks. Each cell is recorded as “Learned/Forgotten”. CPL keeps the forgotten count much smaller across both tasks compared with FT.
+Dataset Yahoo AG DBpedia Num-ds Num-cm C-STANCE MMBench
+FT 50/102 25/10 29/5 71/14 54/41 13/13 115/60
+CPL 50/0 30/1 29/0 87/5 57/7 21/0 116/0
+Appendix K Extending CPL to Out-of-Set Generalization
+The previous experiments focus on the ideal scenario where the Mastered Set $\mathcal{M}$ is fully accessible.
+To evaluate the out-of-set generalization capability of CPL, we design an experiment where both Injection Set and Mastered Set are partitioned into training and evaluation subsets, respectively.
+The mask is computed from the training subset, while the evaluation subset is used to assess learning and forgetting performance.
+In Table 14, “Injection” refers to splitting the Injection Set into training and evaluation subsets, while the Mastered Set remains fully accessible for mask computation.
+Since the Mastered Set is available, CPL still forgets zero question.
+However, due to distribution shifts between the training and evaluation subsets of the Injection Set, the number of learned questions is reduced.
+However, CPL still learn comparable number of questions compared with FT.
+“Mastered” refers to splitting the Mastered Set into training and evaluation subsets, while the Injection Set remains fully accessible for training.
+Since the gradient distribution of evaluation subset can only be partially estimated from the training subset, CPL forgets more questions compared with the ideal scenario.
+Nevertheless, CPL still reduces forgetting by 59.1%–81.7% compared with FT, and incorporating the training subset of the Mastered Set using CPL is more effective than using replay.
+Table 14: Performance evaluation under out-of-set setting.
+Qwen2.5 1.5B is selected for experiment.
+To assess generalization, Injection Set and Mastered Set are partitioned into training and evaluation subsets, respectively.
+This table reports the learning and forgetting performance on unseen evaluation data, demonstrating the out-of-set generalization capability of CPL.
+Method MMLU MedQA ARC-C CSQA
+Learned Forgot Learned Forgot Learned Forgot Learned Forgot
+Injection FT 24 48 34 75 29 71 22 114
+(7.3%) (5.5%) (9.7%) (13.2%) (17.9%) (7.5%) (16.2%) (13.8%)
+CPL 31 0 39 0 29 0 23 0
+(9.5%) (0.0%) (11.1%) (0.0%) (17.9%) (0.0%) (16.9%) (0.0%)
+Mastered FT 178 110 208 143 104 49 86 94
+(26.3%) (26.0%) (29.5%) (50.4%) (32.1%) (10.4%) (31.6%) (22.8%)
+Replay 157 61 124 58 86 14 70 28
+(23.2%) (14.4%) (17.6%) (20.4%) (26.5%) (3.0%) (25.7%) (6.8%)
+CPL 179 46 243 33 116 9 147 18
+(26.4%) (10.9%) (34.5%) (11.6%) (35.8%) (1.9%) (54.0%) (4.4%)
+Appendix L Extending CPL to Cross-Prompt Generalization
+To evaluate CPL inject the specific prompt format or the underlying knowledge, cross-prompt generalization is studided.
+Four prompts as following are designed for each question.
+A question enters the Mastered Set only if all prompts are answered correctly, and enters the Injection Set only if all prompts are answered incorrectly.
+The first and third prompts are used for training, while the second and fourth prompts are used for evaluation.
+Table 15 reports CPL mitigates forgetting under unseen prompt formats while maintaining comparable learning performance.
+template1
+Output exactly one uppercase option label.
+Question:
+{question}
+Options:
+A. {option_a}
+B. {option_b}
+C. {option_c}
+D. {option_d}
+Use only A, B, C, or D. Do not explain your answer.
+template2
+Choose the best answer to the question below.
+Question:
+{question}
+Selections:
+A. {option_a}
+B. {option_b}
+C. {option_c}
+D. {option_d}
+Reply with a single capital letter: A, B, C, or D.
+Do not add any explanation.
+template3
+Read the problem below.
+{question}
+Output exactly one number: 1, 2, 3, or 4.
+Choices:
+1) {option_a}
+2) {option_b}
+3) {option_c}
+4) {option_d}
+No reasoning is required.
+template4
+Consider the following question.
+Question:
+{question}
+Answer set:
+1) {option_a}
+2) {option_b}
+3) {option_c}
+4) {option_d}
+Respond with one digit only: 1, 2, 3, or 4.
+No explanation should appear in your response.
+Table 15: Cross-prompt performance of CPL. CPL substantially reduces forgetting under unseen prompt formats while maintaining comparable learning.
+Method MMLU MedQA ARC-C CSQA
+Learned Forgot Learned Forgot Learned Forgot Learned Forgot
+FT 24 263 47 207 22 375 28 321
+CPL 27 98 48 76 24 88 28 75